### v1

# Matérn Kernels for Tunable Implicit Surface Reconstruction

Abstract
We propose to use the family of Matérn kernels for tunable implicit surface reconstruction, building upon the recent success of kernel methods for 3D reconstruction of oriented point clouds. As we show, both, from a theoretical and practical perspective, Matérn kernels have some appealing properties which make them particularly well suited for surface reconstruction—outperforming state-of-the-art methods based on the arc-cosine kernel while being significantly easier to implement, faster to compute, and scaleable. Being stationary, we demonstrate that the Matérn kernels’ spectrum can be tuned in the same fashion as Fourier feature mappings help coordinate-based MLPs to overcome spectral bias. Moreover, we theoretically analyze Matérn kernel’s connection to Siren networks as well as its relation to previously employed arc-cosine kernels. Finally, based on recently introduced Neural Kernel Fields, we present data-dependent Matérn kernels and conclude that especially the Laplace kernel (being part of the Matérn family) is extremely competitive, performing almost on par with state-of-the-art methods in the noise-free case while having a more than five times shorter training time. [^1] Code available at: https://github.com/mweiherer/matern-surface-reconstruction.
1 Introduction
Recovering the shape of objects from sparse or only partial observations is a challenging task. Formally, let $\Omega\subset\mathbb{R}^{d}$ and $\mathcal{X}=\{x_{1},x_{2},\dots,x_{m}\}\subset\Omega$ be a set of $m$ points in $d$ dimensions which forms, together with associated per-point normals, a data set $\mathcal{D}=\mathcal{X}\times\{n_{1},n_{2},\dots,n_{m}\}\subset\Omega\times\mathbb{R}^{d}$. The goal of surface reconstruction is to recover the objects’ shape from $\mathcal{D}$. The shape of objects may be represented as dense point clouds, polygonal meshes, manifold atlases, voxel grids, or (as the zero-level set of) implicit functions, which is the representation we chose to focus on in this work. Specifically, if $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ denotes an implicit function, such as a signed distance function (SDF), then, from a practical perspective, implicit surface reconstruction aims at finding an optimal solution $\hat{f}$ to the following kernel ridge regression (KRR) problem:
$\hat{f}=\operatorname*{arg\,min}_{f\in H}\left\{\sum_{i=1}^{m}|f(x_{i})|^{2}+\|\nabla f(x_{i})-n_{i}\|^{2}+\lambda\|f\|_{H}^{2}\right\},$ (1)
where the function space $H:=H(\Omega)$ is usually taken to be a reproducing kernel Hilbert space (RKHS) with associated reproducing kernel $k:\Omega^{2}\longrightarrow\mathbb{R}$, and $\lambda>0$. Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (Lorensen & Cline, 1987). It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq. (1) turns into the following constrained minimization problem
$\operatorname*{minimize}_{f\in H}\|f\|_{H}\quad\text{s.t.}\quad f(x_{i})=0\quad\text{and}\quad\nabla f(x_{i})=n_{i}\quad\forall i\in\{1,2,\dots,m\}$ (2)
which makes it easier to see how $\hat{f}$ (thus the estimated surface $\mathcal{\hat{S}}$), behaves away from the input points, $\mathcal{X}$. Clearly, while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing. The properties of $\|f\|_{H}$ in $H$ are uniquely determined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq. (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$. Here, $\alpha\in\mathbb{R}^{m}$ and $K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{m\times m}$. This shows that the behavior of the norm in Eq. (2) and hence, of estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by choosing different kernel functions, effectively allowing us to inject certain inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
((a))
((b))
Figure 1: Matérn kernels with associated spectral densities. We propose to use the family of Matérn kernels for tunable implicit surface reconstruction, parametrized by a smoothness parameter, $\nu>0$, that controls the differentiability of the kernel, and a bandwidth parameter $h>0$. Both parameters allow explicit manipulation of the kernel and its spectrum. Importantly, Matérn kernels recover the Laplace kernel for $\nu=1/2$ and the Gaussian kernel as $\nu\rightarrow\infty$.
Even further, if the chosen kernel has adjustable parameters, they can be used to adaptively (on a shape-by-shape basis) manipulate the inductive bias.
Although already used in the mid-90s and early 2000s (Savchenko et al., 1995; Carr et al., 1997; 2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) eventually evolving as the new state-of-the-art. While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (Williams et al., 2021; 2022) employ the first-order arc-cosine kernel. Introduced by Cho & Saul (2009), arc-cosine kernels have been shown to mimic the computation of (two-layer) fully-connected ReLU networks; indeed, if the layer-width tends to infinity, the first-order arc-cosine kernel is identified with the Neural Tangent Kernel (NTK; Jacot et al. (2018)) of the network.
Figure 2: Tunable Matérn kernels lead to better surface reconstructions compared to the previously employed arc-cosine kernel. Here, we show surface reconstructions from sparse point clouds of just 1,000 points.
One important aspect of this kernel is that it does not have adjustable parameters. While this can be advantageous in some situations, it also prevents users from tuning their reconstructions. Clearly, if the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
In this work, we propose to use a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2. Contrary to arc-cosine kernels, Matérn kernels are stationary (hence rotation- and translation-invariant), locally-supported (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace (for $\nu=1/2$) and Gaussian kernel (as $\nu\rightarrow\infty$). As we will show, both, from a theoretical and practical perspective, Matérn kernels have some intriguing properties which make them particularly well suited for surface reconstruction. Most notably, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details). In addition, our theoretical analysis reveals that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activation function; Sitzmann et al. (2020)) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs. Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (i.e., the Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the hypersphere $\mathbb{S}^{d-1}$.
Our experimental evaluation demonstrates that Matérn 1/2 and 3/2 are extremely competitive, outperforming the arc-cosine kernel while being significantly easier to implement (essentially two lines of standard PyTorch code), faster to compute, and scalable. Moreover, by leveraging Neural Kernel Fields (Williams et al., 2022), we show that data-dependent (a.k.a. learnable) Matérn kernels perform almost on par with highly sophisticated and well-engineered NKSR in the noise-free case while being more than five times faster to train.
2 Related Work
We briefly review some relevant literature about 3D implicit surface reconstruction from oriented point clouds, focusing on kernel-based methods. For a more in-depth overview, including neural-network-based reconstruction methods, see surveys by Berger et al. (2017); Huang et al. (2022b).
Early approaches for kernel-based 3D surface reconstruction mostly employ radial basis functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005). Nowadays, the probably most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; (Kazhdan & Hoppe, 2013)) which, however, can itself be viewed as kernel method as shown in Williams et al. (2021). Only recently, Neural Splines (NS; Williams et al. (2021)) proposed to use the so-called (first-order) arc-cosine kernel,
$k_{\text{AC}}(x,y)=\frac{\|x\|\|y\|}{\pi}\left(\sin\theta+(\pi-\theta)\cos\theta\right),\quad\text{where}\quad\theta=\cos^{-1}\left(\frac{x^{\top}y}{\|x\|\|y\|}\right)$ (3)
for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques as well as non-linear methods based on neural networks by a large margin. This method laid the cornerstone for the Neural Kernel Field (NKF; Williams et al. (2022)) framework which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016). Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn a set of spatially-varying basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense. Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose surface reconstruction method that is highly scaleable and robust against noise, eventually becoming state-of-the-art.
3 Background: Matérn Kernels
We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) instead of the arc-cosine kernel for implicit surface reconstruction, being defined as:
$k_{\nu}(x,y)=\Phi_{\nu}(\|x-y\|)=\Phi_{\nu}(\tau)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\tau}{h}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\tau}{h}\right),$ (4)
where $\nu>0$ is a smoothness parameter that explicitly controls the differentiability, and $h>0$ is known as the shape parameter (or bandwidth) of the kernel. $\Gamma$ denotes the Gamma function, and $K_{\nu}$ is the modified Bessel function of the second kind of order $\nu$. For $\nu=p+\frac{1}{2}$, $p\in\mathbb{N}^{+}$ (i.e., if $\nu$ is a half-integer), Eq. (4) simplifies to a product of an exponential and polynomial of degree $p$, that is
$\Phi_{p+\frac{1}{2}}(\tau)=\exp\left(-\frac{\sqrt{2p+1}\tau}{h}\right)\frac{p!}{(2p)!}\sum_{i=0}^{p}\frac{(p+i)!}{i!(p-i)!}\left(\frac{2\sqrt{2p+1}\tau}{h}\right)^{p-i}.$ (5)
Matérn kernels generalize a variety of other kernels; using Eq. (5), the most popular ones can be written in closed-form as:
$p$ $=0:\Phi_{\text{1/2}}(\tau)=\exp\left(-\frac{\tau}{h}\right),$ (6)
$p$ $=1:\Phi_{\text{3/2}}(\tau)=\exp\left(-\frac{\sqrt{3}\tau}{h}\right)\left(1+\frac{\sqrt{3}\tau}{h}\right),$ (7)
$p$ $=2:\Phi_{\text{5/2}}(\tau)=\exp\left(-\frac{\sqrt{5}\tau}{h}\right)\left(1+\frac{\sqrt{5}\tau}{h}+\frac{5\tau^{2}}{3h^{2}}\right),$ (8)
where $\Phi_{\text{1/2}}$ is known as the Laplace kernel. In the limiting case of $\nu\rightarrow\infty$, Matérn kernels recover the popular Gaussian kernel:
$\Phi_{\infty}(\tau):=\lim_{\nu\rightarrow\infty}\Phi_{\nu}(\tau)=\exp\left(-\frac{\tau^{2}}{2h^{2}}\right).$ (9)
The spectral density of Matérn kernels is given as the inverse Fourier transform of $\Phi_{\nu}$, that is
$p_{\nu}(\omega)=\mathcal{F}^{-1}[\Phi_{\nu}]=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$ (10)
which is due to the fact that Matérn kernels are stationary (or shift-invariant—they only depend on the difference $x-y$ between two points, $x,y\in\Omega$); hence, kernel and associated spectral density are Fourier duals (a consequence of Bochner’s Theorem). Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$.
The RKHS induced by Matérn kernels is quite well understood (see, e.g., (Kanagawa et al., 2018)) and can be formulated as $H_{\nu}=\{f\in L_{2}(\mathbb{R}^{d})\cap C(\mathbb{R}^{d}):\|f\|_{H_{\nu}}^{2}<\infty\}$, where the RKHS norm is
$\|f\|_{H_{\nu}}^{2}=h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}+\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega.$ (11)
For more information about Matérn kernels, we refer the interested reader to Porcu et al. (2023) for a recent review. Next, we revisit some basic properties of Matérn kernels.
3.1 Basic Properties
Differentiable. Matérn kernels allow for controlled smoothness, being exactly $\lceil\nu\rceil-1$ times differentiable (in the mean-square sense); for instance, while the Matérn 1/2 kernel is zero times differentiable, the Gaussian kernel (for $\nu\rightarrow\infty$) is infinitely often differentiable.
Stationary. Matérn kernels are stationary, i.e., they only depend on the difference $x-y$ between two points $x,y\in\Omega$. If the distance is Euclidean, Matérn kernels are also isotropic (or homogeneous); they only depend on $\|x-y\|$. As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of $\mathcal{X}$. We note that, while the previously used arc-cosine kernel is also rotation invariant, it is not translation invariant.
Locally supported. As opposed to arc-cosine kernels, Matérn kernels are locally supported (or compact); that is, they tend to zero as the distance $\|x-y\|$ becomes large. This is a very desirable property, especially from a computational perspective, as it leads to sparse Gram matrices when multiplied with the kernel $\Phi^{\prime}_{\nu^{\prime}}(\tau)=\max\left\{0,\left(1-\tau/h^{\prime}\right)^{\nu^{\prime}}\right\}$ for suitable $\nu^{\prime}$ and $h^{\prime}$, effectively allowing the use of highly efficient and scalable sparse linear solvers. Please see Genton (2001) on how to choose suitable values for $\nu^{\prime}$ and the cut-off distance, $h^{\prime}$.
((a))
((b))
Figure 3: Eigenvalue decay of Matérn kernels. While we fix $h=1$ and vary $\nu$ on the left, the EDR for $\nu=1/2$ and different values of $h$ is shown on the right. Larger values of $\nu$ and $h$, i.e., smoother kernels, lead to faster eigenvalue decay; hence slow convergence to high-frequency details.
4 Matérn Kernels for Tunable Surface Reconstruction
We now derive new theoretical insights into Matérn kernels, ultimately aiming to provide arguments as to why we believe this family of kernels is particularly well suited for surface reconstruction.
4.1 Eigenvalue Decay, Spectral Bias, and Reconstruction Error
Eigenvalue decay and spectral bias. We begin by analyzing Matérn kernel’s eigenvalue decay rate (EDR) and show that it can be tuned by varying the smoothness parameter $\nu$ and/or bandwidth $h$. As investigated in Tancik et al. (2020) for coordinate-based MLPs, a rapid decrease of the associated NTK’s eigenvalues implies slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components). Consequently, a smaller EDR, i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content; hence, in the context of implicit surface reconstruction, more detailed geometry. To overcome this slow convergence, a phenomenon known as spectral bias, Tancik et al. (2020) propose to use a (random) Fourier feature mapping of the form $\gamma(x)=[\cos(2\pi Bx),\sin(2\pi Bx)]$ applied to the inputs $x\in\mathbb{R}^{d}$ before passing them to the MLP, effectively transforming the MLP’s NTK, $k_{\text{NTK}}$, into a stationary kernel with tunable bandwidth parameter, $k_{\text{NTK}}(\gamma(x)^{\top}\gamma(y))=k_{\text{NTK}}(k(x-y))=:k^{\prime}_{\text{NTK}}(x-y)$. When $B\in\mathbb{R}^{q\times d}$ is drawn from a standard normal distribution (i.e., zero mean and unit variance), $k$ is an approximation of the Gaussian kernel (Rahimi & Recht, 2007). This allows explicit manipulation of the kernels’ spectrum, overcoming spectral bias.
Matérn kernels, being stationary, allow for the same degree of control over the EDR as a Fourier feature mapping; their spectrum can also be tuned (by varying $\nu$ and/or $h$, see Figure 1). This is in stark contrast to previously employed, non-stationary and parameter-less arc-cosine kernels, whose spectrum is not tunable. To further study the dependence of the EDR on $\nu$ and $h$, we use:
Theorem 1 (Seeger et al. (2008), Theorem 3).
The eigenvalues of Matérn kernels decay polynomially at rate
$\Theta\left(h^{-2\nu}s^{-(1+2\nu/d)}\right)$
with bandwidth parameter $h>0$, and for finite smoothness $0<\nu<\infty$.
Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by Li et al. (2024) for general input domains and distributions. As seen, the smoother the kernel (i.e., the higher $\nu$) and the higher $h$, the faster the eigenvalues decay. This implies less detailed surface reconstruction (with a higher error) for smoother kernels and greater $h$; indeed, that is exactly what we observe in practice, see Figure 4. Contrary, by lowering both parameters, one can achieve faster convergence to high-frequency geometric details, effectively leading to more nuanced surface reconstruction and a smaller reconstruction error. Interestingly, the first-order arc-cosine kernel and Matérn 1/2 have the same EDR if $h=1$.
Figure 4: Matérn kernels are tunable. Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3). However, setting $h$ too small ($<0.5$) results in over-fitting, while setting $h$ too big ($>10$) over-smooths (i.e., under-fits) the true surface. This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
Reconstruction error. We proceed by investigating the bandwidth’s influence on the reconstruction error, measured by the mean-squared error which is defined (and can be bounded) as:
$\|f-\hat{f}\|_{L_{2}}=\left(\int_{\Omega}\left(f(x)-\hat{f}(x)\right)^{2}\mathrm{d}x\right)^{1/2}\leq C_{\mathcal{X},\Omega}^{(\nu+d)/2}\|f\|_{H_{\nu}},$ (12)
where $C_{\mathcal{X},\Omega}$ is a constant that only depends on $\mathcal{X}$ and $\Omega$ and for $\nu<\infty$. For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., Santin & Schaback (2016). By inspecting Eq. (12), we notice that the magnitude of the RKHS norm has a considerably high influence on the reconstruction error. Based on this observation and the fact that $\|f\|_{H_{\nu}}$ depends on $h$, our goal is now to further study the effect of $h$ on the RKHS norm. As a first result, we have:
Proposition 2.
The RKHS norm of Matérn kernels as defined in Eq. (11) can be bounded by
$\|f\|^{2}_{H_{\nu}}\leq h^{2\nu}\left(\frac{1}{h^{2\nu+d}}C_{d,\nu}^{1}(\mathcal{F}[f])+C^{2}_{d,\nu}(\mathcal{F}[f])\right),$
where $h>0$, and $C^{1}_{d,\nu}$ and $C^{2}_{d,\nu}$ are functions of $\mathcal{F}[f]$ that do not depend on $h$.
A proof can be found in Appendix A. We immediately see that $\|f\|^{2}_{H_{\nu}}\rightarrow\infty$ in either cases, $h\rightarrow 0$ and $h\rightarrow\infty$. Moreover, based on Proposition 2, it is easy to show that the norm, as a function of $h$, decreases as $h\rightarrow h^{*}$ from the left, and increases as $h\rightarrow\infty$, where $h^{*}$ is given as
$h^{*}=\left(\frac{d}{2\nu}Q\right)^{1/(2\nu+d)}\quad\text{with}\quad Q=\frac{C^{1}_{d,\nu}(\mathcal{F}[f])}{C^{2}_{d,\nu}(\mathcal{F}[f])}$ (13)
(see Appendix B for a derivation). This can also be seen from the inset figure on the right, where we plot the RKHS norm with respect to the obtained bound as a function of $h$.
Interestingly, our analysis reveals that the reconstruction error can not be made arbitrarily small by just lowering $h$; if the bandwidth $h$ is chosen too small, the reconstruction error starts to increase again. Conversely, if $h$ is set too high, the resulting surface is too smooth, ultimately leading to increased errors. In fact, what we actually observe is that Matérn kernels under-fit the true surface when $h$ is set too big, and, conversely, over-fit if $h$ is chosen too small, see also Figure 4. As we have shown, the optimal trade-off is reached if $h$ equals $h^{*}$. Moreover, whilst the bandwidth $h$ does play a critical role, our results show it is not very sensitive in practice, especially for Matérn 1/2 and 3/2 (see inset figure).
4.2 Relation to Neural Networks and the Arc-cosine Kernel
Relation to neural networks. Next, we study the connection between Matérn kernels and neural networks. It is well known that first-order arc-cosine kernels mimic the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al. (2021). Here, $[x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens Sitzmann et al. (2020) (fully-connected networks with sine activation function) if the layer width approaches infinity. We claim the following:
Theorem 3.
Consider a two-layer fully-connected network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ with sine activation function, $m$ hidden neurons, and fixed bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$. Let $h,\nu>0$ be given. If $w_{i}$ is randomly initialized from
$p_{\nu}(\omega)=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$
and $b_{i}\sim\mathcal{U}(0,2\pi)$, the NTK of $f$ is a Matérn kernel with bandwidth $h$ and smoothness $\nu$ as $m\rightarrow\infty$.
A proof can be found in Appendix C. This result establishes for the first time a connection between Matérn kernels and the computation in neural networks, in particular, Sirens.
Matérn kernels vs. arc-cosine kernels. Lastly, we aim to compare Matérn kernels with the previously used first-order arc-cosine kernels. Our analysis is based on:
Theorem 4 (Chen & Xu (2021), Theorem 1).
When restricted to the hypersphere $\mathbb{S}^{d-1}$, the RKHS of the Matérn kernel with smoothness $\nu=1/2$ (the Laplace kernel) include the same set of functions as the RKHS induced by the NTK of a fully-connected ($L\geq 2$)-layer ReLU network.
In other words, Theorem 4 shows that the RKHS of the Laplace kernel (a Matérn kernel with smoothness $\nu=1/2$) is the same as the RKHS induced by the NTK of a fully-connected ReLU network when inputs are restricted to $\mathbb{S}^{d-1}$. This immediately raises the question of whether the NTK for general input domains is significantly different from a standard Laplace kernel. Indeed, empirical results (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020) already suggest a negative answer; however, from a theoretical perspective, the claim is still unproven. Based on Theorem 4, we obtain the following connection between Laplace and the first-order arc-cosine kernel:
Corollary 5.
When restricted to the hypersphere $\mathbb{S}^{d-1}$, (1) the RKHS of the Matérn kernel with smoothness $\nu=1/2$ (the Laplace kernel) is equal to the RKHS induced by the first-order arc-cosine kernel, implying that (2) the Laplace and arc-cosine kernel are equal up to an affine transformation.
See Appendix D for a proof. For general input domains, our empirical results presented next support the claim mentioned above (at least for ($L=2$)-layer networks), showing that the Laplace kernel even outperforms the arc-cosine kernel.
5 Experiments and Results
We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (Chang et al., 2015) and the Surface Reconstruction Benchmark (SRB; Berger et al. (2013)) in Section 5.1. For ShapeNet, we use the train/val/test split provided by Mescheder et al. (2019). Leveraging Neural Kernel Fields (Williams et al., 2022), we present and evaluate learnable Matérn kernels in Section 5.2.
5.1 Surface Reconstruction on ShapeNet and SRB
ShapeNet. We compare Matérn kernels in a sparse setting against the arc-cosine kernel as implemented in Neural Splines (NS, Williams et al. (2021)). To do so, we randomly sample $m=1,000$ surface points with corresponding normals for each shape. We implemented Matérn kernels in PyTorch and took the official CUDA implementation of the arc-cosine kernel from NS, eventually integrated into a unified framework to ensure fair comparison. Notably, we did not use Nyström sampling as in NS, and employ PyTorch’s direct Cholesky solver to solve the KRR problem in Eq. (1) instead of an iterative conjugate gradient solver. Moreover, similar to NS, we approximate the gradient part in Eq. (1) with finite differences. See Appendix E for details.
F-Score $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 Matérn 1/2 93.6 93.7 93.7 4.05 4.02 4.02 Matérn 3/2 94.6 94.8 94.9 4.05 4.00 4.09 Matérn 5/2 92.4 93.8 92.9 6.42 5.65 6.91 Matérn $\infty$ 59.3 52.4 40.9 57.56 50.59 48.38 Arc-cosine 92.8 4.67 NS 90.6 4.74 SPSR 84.3 6.26       CD $\downarrow$ HD $\downarrow$ Time Matérn 1/2 0.21 4.43 9.85 sec Matérn 3/2 0.18 2.93 12.47 sec Matérn 5/2 0.58 22.52 15.26 sec Matérn $\infty$ 3.83 33.27 12.76 sec NS* 0.19 3.19 19.89 sec NS 0.17 2.85 11.91 sec SPSR 0.21 4.69 1.65 sec FFN 0.28 4.45 337.91 sec Siren 0.19 3.86 599.00 sec
[…]
Quantitative results in terms of F-Score (with a threshold of 0.01) and Chamfer distance (CD; always reported $\times$ 103) can be found in Table 1. Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (Kazhdan & Hoppe, 2013). In addition, our experiments show that the bandwidth parameter, $h$, can be used to tune the surface reconstruction (lower the error), while not being too sensitive in practice. Matérn 5/2 and the Gaussian kernel (for $\nu\rightarrow\infty$) perform significantly worse, being simply too smooth. Figure 2 and Appendix E shows qualitative results.
SRB. Additionally, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark (Berger et al., 2013) which consists of five complex shapes simulated from noisy range scans. For this, we integrated Matérn kernels into the NS framework, using KeOps (Charlier et al., 2021) and the FALKON (Rudi et al., 2017) solver. We used the same parameter settings as in the original NS paper. In addition, for Matérn kernels, we did a modest parameter sweep over the bandwidth, $h\in\{0.5,1,2\}$, and took the reconstruction with the lowest Chamfer distance.
Figure 5: Qualitative results for non-learnable kernels on SRB (left) and learned kernels on ShapeNet (right). Please see Appendix F and G for more qualitative results.
Quantitative results obtained by employing Chamfer and Hausdorff distance (HD) can be found in Table 1, demonstrating that Matérn 3/2 performs on par with NS while being faster to compute (compared to what we measured for NS). See Figure 5 (left) and Appendix F for qualitative results and per-object metrics.
5.2 Data-dependent Matérn Kernels
We leverage the Neural Kernel Field (NKF) framework introduced by Williams et al. (2022) to make Matérn kernels learnable. Specifically, NKF proposed to feed points $x,y\in\mathbb{R}^{d}$ through an input-conditioned neural network $\varphi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{q}$ before evaluating a kernel function, $k^{\varphi}(x,y;\mathcal{D})=k\left([x,\varphi(x;\mathcal{D})],[y,\varphi(y;\mathcal{D})]\right)$. We re-implemented the NKF framework since the authors did not provide code. We set $q=32$ for our experiments. Please refer to (Peng et al., 2020; Williams et al., 2022) for more details.
Sparse surface reconstruction and extreme generalization. We compare learned Matérn kernels against the arc-cosine kernel originally used within the NKF framework and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet. Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels. Additionally, we evaluate Matérn kernel’s out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories. Table 2 reports the results, quantified using intersection-over-union (IoU), Chamfer distance, and normal consistency (NC). Learned Matérn 1/2 and 3/2 outperform the arc-cosine kernel while being significantly faster to train. Training NKF on the entire ShapeNet dataset (consisting of approx. 30,000 shapes) takes about six days—almost twice as long as for Matérn kernels, which require about three days (measured on a single NVIDIA A100). Additionally, although Matérn 1/2 is not able to surpass NKSR, it comes very close (95.3 vs. 95.6 in IoU) while having a shorter training time. Regarding out-of-category generalization, we observe that Matérn 1/2 performs best, followed by Matérn 3/2, NKF, and NKSR. Please see Figure 5 (right) and Appendix G for qualitative results.
IoU $\uparrow$ CD $\downarrow$ NC $\uparrow$ Matérn 1/2 95.3 2.58 95.6 Matérn 3/2 94.9 2.70 95.3 Matérn 5/2 93.3 3.07 94.9 Matérn $\infty$ 92.1 3.39 94.2 NKF* 94.7 2.70 95.2 NKF 94.7 2.65 95.4 NKSR 95.6 2.34 95.4       Train only on chairs, test on all IoU $\uparrow$ CD $\downarrow$ NC $\uparrow$ Time/epoch Matérn 1/2 93.4 3.06 94.3 7.71 min Matérn 3/2 92.8 3.34 94.2 8.26 min Matérn 5/2 90.5 4.11 93.9 8.56 min Matérn $\infty$ 84.7 6.70 92.6 7.74 min NKF* 92.8 3.30 94.1 31.98 min NKSR 89.6 2.70 94.1 41.66 min
[…]
No noise ($\sigma=0$) Small noise ($\sigma=0.0025$) Big noise ($\sigma=0.005$) IoU $\uparrow$ CD $\downarrow$ IoU $\uparrow$ CD $\downarrow$ IoU $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 0.5 1 2 0.5 1 2 0.5 1 2 0.5 1 2 Matérn 1/2 93.3 93.6 93.5 2.88 2.85 2.87 91.9 92.1 92.1 3.03 3.04 3.07 89.5 89.6 89.6 3.37 3.36 3.39 Matérn 3/2 92.7 92.7 92.3 3.08 3.20 3.25 89.7 89.8 89.7 3.48 3.69 3.54 85.8 85.5 86.3 4.23 4.32 4.15 Matérn 5/2 91.8 90.5 88.8 3.58 3.69 4.18 87.9 87.6 87.0 4.29 4.48 4.43 83.4 82.2 84.3 4.90 5.03 5.11 Matérn $\infty$ 86.4 85.8 86.8 5.80 5.07 4.75 81.3 83.3 81.0 7.27 5.62 6.45 76.7 80.7 79.3 8.27 5.90 6.67 NKF* 93.2 2.97 92.0 3.12 89.5 3.39 NKSR 91.1 2.65 90.1 2.98 88.4 3.41
Table 3: Robustness against noise. We compare Matérn kernels’ robustness against noise on a subset of the ShapeNet dataset (including about 1,700 shapes) using different noise levels, $\sigma\in\{0,0.0025,0.005\}$ and bandwidths, $h\in\{0.5,1,2\}$. Again, NKF* denotes a re-implementation of NKF. Bold marks best result, underline second best.
Robustness against noise. We evaluate Matérn kernel’s robustness against different noise levels, $\sigma\in\{0,0.0025,0.005\}$, on a subset of the ShapeNet dataset which includes approximately 1,700 shapes. To construct the dataset, we downsampled each ShapeNet category to include only 5% of the shapes. Table 3 presents the results. NKSR, being specifically optimized to deal with noisy inputs, degrades the least with increasing noise level (3% in IoU from no to big noise versus 4.3% for Matérn 1/2). Moreover, Matérn 1/2 is slightly more robust against noise than NKF. Generally, varying the bandwidth, $h$, helps increase the robustness against noise.
Convergence speed. Finally, we analyze Matérn kernel’s convergence speed in comparison to the arc-cosine kernel as implemented in NKF, see inset figure on the right. As exemplified, Matérn 1/2 already converges after just 100 epochs, while the arc-cosine kernel requires twice as many epochs, also never being able to reach the same top accuracy. In general, we see that the smoother the kernel, the slower the convergence—an observation that perfectly matches our theoretical analysis in Section 4.1. Matérn 1/2 converges the fastest, followed by Matérn 3/2, 5/2, and the Gaussian kernel.
6 Conclusion
In this work, we have proposed to use the family of Matérn kernels for tunable implicit surface reconstruction, acting as a drop-in replacement of the recently employed first-order arc-cosine kernel within frameworks such as Neural Splines (Williams et al., 2021) and Neural Kernel Fields (Williams et al., 2022). We presented an in-depth theoretical analysis of Matérn kernels, analyzing its relation to Fourier feature mappings, Siren networks, as well as arc-cosine kernels. Our systematic evaluation confirmed our theoretical results and showed that Matérn kernels are significantly faster to compute and train, and, especially Matérn 1/2 and 3/2, lead to better surface reconstructions compared to the arc-cosine kernel—both, in a non-learnable as well as learnable regime. In the future, we plan to incorporate Matérn kernels into the NKSR framework to enhance robustness against noise.
Acknowledgments
We thank Francis Williams for valuable discussions and his support in re-implementing the NKF framework and Lukas Meyer for his help during NKSR evaluation. This work was funded by the German Federal Ministry of Education and Research (BMBF), FKZ: 01IS22082 (IRRW). The authors are responsible for the content of this publication.
Appendix A Proof of Proposition 2
See 2
Proof.
Based on a simple bound
$I$ $=\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}+\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega$
$\leq\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega+\int_{\mathbb{R}^{d}}\left(\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega$
$=\left(\frac{2\nu}{h^{2}}\right)^{\nu+d/2}\underbrace{\int_{\mathbb{R}^{d}}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega}_{=:C^{\prime}(\mathcal{F}[f])}+\underbrace{\int_{\mathbb{R}^{d}}\left(\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega}_{=:C^{\prime}_{d,\nu}(\mathcal{F}[f])}$
$=\left(\frac{2\nu}{h^{2}}\right)^{\nu+d/2}C^{\prime}(\mathcal{F}[f])+C^{\prime}_{d,\nu}(\mathcal{F}[f])$
we get
$\|f\|_{H_{\nu}}^{2}$ $=h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}I$
$\leq h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}\left(\left(\frac{2\nu}{h^{2}}\right)^{\nu+d/2}C^{\prime}(\mathcal{F}[f])+C^{\prime}_{d,\nu}(\mathcal{F}[f])\right)$
$=h^{2\nu}\left(\frac{1}{h^{2\nu+d}}\underbrace{(2\nu)^{\nu+d/2}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}C^{\prime}(\mathcal{F}[f])}_{=:C^{1}_{d,\nu}(\mathcal{F}[f])}+\underbrace{\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}C^{\prime}_{d,\nu}(\mathcal{F}[f])}_{=:C^{2}_{d,\nu}(\mathcal{F}[f])}\right)$
$=h^{2\nu}\left(\frac{1}{h^{2\nu+d}}C^{1}_{d,\nu}(\mathcal{F}[f])+C^{2}_{d,\nu}(\mathcal{F}[f])\right)$
which concludes the proof. ∎
Appendix B Derivation of $h^{*}$ in Eq. (13)
Taking derivative of the bound presented in Proposition 2 w.r.t. $h$ yields
$\frac{\mathrm{d}}{\mathrm{d}h}\left(h^{2\nu}\left(\frac{1}{h^{2\nu+d}}C^{1}_{d,\nu}(\mathcal{F}[f])+C^{2}_{d,\nu}(\mathcal{F}[f])\right)\right)=\frac{1}{h^{d+1}}\left(2\nu C^{2}_{d,\nu}(\mathcal{F}[f])h^{2\nu+d}-dC^{1}_{d,\nu}(\mathcal{F}[f])\right).$
Since $d,h>0$, the leading factor $1/h^{d+1}$ never becomes zero, so we must have
$2\nu C^{2}_{d,\nu}(\mathcal{F}[f])h^{2\nu+d}-dC^{1}_{d,\nu}(\mathcal{F}[f])=0\iff h^{*}=\left(\frac{d}{2\nu}\frac{C^{1}_{d,\nu}(\mathcal{F}[f])}{C^{2}_{d,\nu}(\mathcal{F}[f])}\right)^{1/(2\nu+d)}$
which shows what we have stated in Eq. (13) of the main paper.
Appendix C Proof of Theorem 3
The proof of Theorem 3 is based on Bochner’s theorem (in harmonic analysis) which reads:
Theorem 6 (Bochner).
A continuous function $k:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ with $k(0)=1$ is positive definite if and only if there exists a finite positive Borel measure $\mu$ on $\mathbb{R}^{d}$ such that
$k(\tau)=\int_{\mathbb{R}^{d}}e^{i\omega^{\top}\tau}\mathrm{d}\mu(\omega).$
In essence, Bochner’s theorem states that $k$ and $\mu$ are Fourier duals; given a stationary kernel $k$, one can obtain $\mu$ (better known as spectral density if normalized) by applying the inverse Fourier transform to $k$. On the other hand, given a spectral density $\mu$, one can obtain the corresponding kernel function $k$ by applying the Fourier transform to $\mu$. Based on Bochner’s theorem,
$k(\tau)=k(x-y)=\int_{\mathbb{R}^{d}}e^{i\omega^{\top}(x-y)}\mathrm{d}\mu(\omega)=\mathbb{E}_{\omega\sim\mu}\left[\cos(\omega^{\top}(x-y))\right]$
because $k$ is real and for even spectral densities. On the other hand, Rahimi & Recht (2007) showed
$k(\tau)=k(x-y)$ $=\mathbb{E}_{\omega\sim\mu}[\cos(\omega^{\top}(x-y))]$
$=\mathbb{E}_{\omega\sim\mu}[\cos(\omega^{\top}(x-y))]+\underbrace{\mathbb{E}_{\omega\sim\mu,b\sim\mathcal{U}(0,2\pi)}[\cos(\omega^{\top}(x+y)+2b)]}_{=0\,(*)}$
$=\mathbb{E}_{\omega\sim\mu,b\sim\mathcal{U}(0,2\pi)}[\cos(\omega^{\top}(x-y))+\cos(\omega^{\top}(x+y)+2b)]$
$=\mathbb{E}_{\omega\sim\mu,b\sim\mathcal{U}(0,2\pi)}[2\cos(\omega^{\top}x+b)\cos(\omega^{\top}y+b)],$
where ()(*)( ∗ ) follows because $b$ is uniformly distributed in $[0,2\pi]$ (hence, the inner expectation w.r.t. $b$ becomes zero).
We will now prove Theorem 3, copied below for convenience:
See 3
Proof.
Let $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ be a two-layer fully-connected network with sine activation function, i.e.,
$f(x)=\sqrt{\frac{2}{m}}\sum_{i=1}^{m}v_{i}\sin(w_{i}^{\top}x+b_{i})$
with bottom-layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$, biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$, and top-layer weights $v=(v_{1},v_{2},\dots,v_{m})\in\mathbb{R}^{m}$. Assume that $W$ and $b$ are fixed (that is, we only allow the second layer to be trained), and initialized according to some distribution. Choose
$w_{i}\sim p_{\nu}(\omega)=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$
for $h,\nu>0$ and $b_{i}\sim\mathcal{U}(0,2\pi)$ for all $i\in\{1,2,\dots,n\}$. Then, based on the fact that $\partial_{v_{i}}f(x)=\sqrt{2}\sin(\omega_{i}^{\top}x+b_{i})/\sqrt{m}$, it is easy to see that the NTK of $f$ is given by
$k_{\text{NTK}}(x,y)$ $=\sum_{i=1}^{m}\partial_{v_{i}}f(x)\partial_{v_{i}}f(y)$
$=\frac{1}{m}\sum_{i=1}^{m}2\sin(w_{i}^{\top}x+b_{i})\sin(w_{i}^{\top}y+b_{i})$
$=\frac{1}{m}\sum_{i=1}^{m}2\cos(w_{i}^{\top}x+b_{i})\cos(w_{i}^{\top}y+b_{i})-2\cos(w_{i}^{\top}(x+y)+2b_{i})$
$=\frac{1}{m}\sum_{i=1}^{m}2\cos(w_{i}^{\top}x+b_{i})\cos(w_{i}^{\top}y+b_{i})-\frac{1}{m}\sum_{i=1}^{m}2\cos(w_{i}^{\top}(x+y)+2b_{i}).$
As $m\rightarrow\infty$, we finally obtain
$k_{\text{NTK}}(x,y)$ $=\mathbb{E}_{(w,b)}\left[2\cos(w^{\top}x+b)\cos(w^{\top}y+b)\right]-2\underbrace{\mathbb{E}_{(w,b)}\left[\cos(w^{\top}(x+y)+2b)\right]}_{=0\,(*)}$
${=}k_{\nu}(x,y)$
where ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see Eq. (10) of the main paper. ∎
Appendix D Proof of Corollary 5
We first re-state the following result, needed to prove the second part of Corollary 5:
Theorem 7 (Saitoh & Sawano (2016), Theorem 2.17).
Let $k_{1}$ and $k_{2}$ be two positive definite kernels. Denote by $H_{1},H_{2}$ the RKHSs induced by $k_{1}$ and $k_{2}$. Then,
$H_{1}\subset H_{2}\iff\gamma^{2}k_{2}-k_{1}\text{ is positive definite for }\gamma>0.$
We are now ready to prove Corollary 5 which is copied below for convenience:
See 5
Proof.
(1) Follows from Theorem 4 by noting that the NTK of a fully-connected $(L=2$)-layer ReLU network is the first-order arc-cosine kernel, see, e.g.Cho & Saul (2009); Williams et al. (2021). For (2), we know that $H_{\nu}\subset H_{\text{AC}}$ and $H_{\text{AC}}\subset H_{\nu}$, so, based on Theorem 7, $k:=\gamma^{2}k_{\text{AC}}-k_{\nu}$ and $\bar{k}:=\bar{\gamma}^{2}k_{\nu}-k_{\text{AC}}$ are both valid kernels for $\gamma,\bar{\gamma}>0$. Re-arranging both equations, we have
$k_{\text{AC}}=ak_{\nu}+b,\quad\text{where}\quad a=1/\gamma^{2},b=k/\gamma^{2}$
and, equivalently,
$k_{\nu}=\bar{a}k_{\text{AC}}+\bar{b},\quad\text{where}\quad\bar{a}=1/\bar{\gamma}^{2},\bar{b}=\bar{k}/\bar{\gamma}^{2}.$
This shows that the Laplace and arc-cosine kernel are indeed related by an affine transformation when restricted to $\mathbb{S}^{d-1}$, concluding the proof. ∎
Appendix E Additional Details for ShapeNet Experiment
In this section, we provide additional details on the ShapeNet evaluation as presented in Section 5.1 of the main paper.
E.1 Implementation Details
Following common practice (see, e.g., Williams et al. (2022)), we use finite differences to approximate the gradient part of the KRR problem in Eq. (1) of the main paper. Specifically, denote $\mathcal{X}^{\prime}:=\mathcal{X}^{+}\cup\mathcal{X}^{-}$, where $\mathcal{X}^{+}:=\{x_{1}+{\epsilon}n_{1},x_{2}+{\epsilon}n_{2},\dots,x_{m}+{\epsilon}n_{m}\}$ and $\mathcal{X}^{-}:=\{x_{1}-{\epsilon}n_{1},x_{2}-{\epsilon}n_{2},\dots,x_{m}-{\epsilon}n_{m}\}$ for a fixed ${\epsilon}>0$. The Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) tells us that the solution to Eq. (1) is of the form
$f(x)=\sum_{x^{\prime}\in\mathcal{X}^{\prime}}\alpha_{i}k(x,x^{\prime})$
which is linear in the coefficients $\alpha\in\mathbb{R}^{2m}$, given by
$\alpha=(K+\lambda I)^{-1}y,\quad\text{where}\quad K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{2m\times 2m},$
and
$y_{i}=\begin{cases}+{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{+},\\ -{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{-}.\end{cases}$
$I\in\mathbb{R}^{2m\times 2m}$ denotes the identity matrix. We employ PyTorch’s built-in Cholesky solver to numerically stable solve for $\alpha$ and set ${\epsilon}=0.005$ for all experiments.
E.2 Further Qualitative Results
We show additional qualitative results on ShapeNet in Figure 6.
Figure 6: More qualitative results on ShapeNet, comparing Matérn 1/2 and 3/2 for different values of $h$ to the arc-cosine kernel as implemented in Neural Splines (Williams et al., 2021).
Appendix F Further Details on the Surface Reconstruction Benchmark
F.1 Experimental Setting
For Neural Splines (NS; Williams et al. (2021)), we used the official implementation provided here: https://github.com/fwilliams/neural-splines. Similar to the original paper, we used 15,000 Nyström samples and did the same parameter sweep over the regularization parameter, $\lambda\in\{0,10^{-13},10^{-12},10^{-11},10^{-10}\}$. For the rest of the parameters (not mentioned in the paper), default values provided in the repository have been used, except for the grid size which we set to 512. We used exactly the same setting for Matérn kernels, except that we did a modest parameter sweep over the bandwidth parameter, $h\in\{0.5,1,2\}$. We utilized the SRB data from here: https://github.com/fwilliams/deep-geometric-prior.
Despite our best efforts, we were unable to reproduce the results for the NS kernel reported in the original paper. We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines) and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance. In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; the authors of (Williams et al., 2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively. In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
To measure runtime, we again used the same setting as in the original NS paper: 15,000 Nyström samples, and $\lambda=10^{-11}$. Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters. Runtime was measured on a single NVIDIA V100, similar to (Williams et al., 2021).
F.2 Additional Results
Per-object metrics are shown in Table 4, and more qualitative result in Figure 7.
Figure 7: Qualitative results on the Surface Reconstruction Benchmark, comparing Matérn 1/2 and 3/2 to Neural Splines (NS; Williams et al. (2021)).
Anchor Daratech DC Gargoyle Lord Quas
CD $\downarrow$ HD $\downarrow$ CD $\downarrow$ HD $\downarrow$ CD $\downarrow$ HD $\downarrow$ CD $\downarrow$ HD $\downarrow$ CD $\downarrow$ HD $\downarrow$
Matérn 1/2 0.31 6.97 0.25 6.11 0.18 2.62 0.19 5.14 0.12 1.32
Matérn 3/2 0.25 5.08 0.23 4.90 0.15 1.24 0.16 2.54 0.12 0.90
Matérn 5/2 0.93 28.33 0.87 32.17 0.34 15.98 0.55 22.52 0.21 9.49
Matérn $\infty$ 2.88 29.48 5.80 45.94 0.35 35.70 3.84 30.70 3.13 24.55
NS 0.27 5.38 0.23 4.67 0.15 1.41 0.17 3.49 0.12 0.99
NS 0.22 4.65 0.21 4.35 0.14 1.35 0.16 3.20 0.12 0.69
SPSR 0.33 7.62 0.26 6.62 0.17 2.79 0.18 4.60 0.12 1.83
FFN 0.31 4.49 0.34 5.97 0.20 2.87 0.22 5.04 0.35 3.90
Siren 0.32 8.19 0.21 4.30 0.15 2.18 0.17 4.64 0.17 0.82
Table 4: Per-object quantitative results on the Surface Reconstruction Benchmark.
Appendix G Additional Results for Data-dependent Kernels
We show additional qualitative results for learned kernels on ShapeNet in Figure 8.
Figure 8: Additional qualitative results for learned kernels on ShapeNet. We compare against Neural Kernel Fields (NKFs; Williams et al. (2022)) and Neural Kernel Surface Reconstruction (NKSR; Huang et al. (2023)).

### v2

@@
 Abstract
-We propose to use the family of Matérn kernels for tunable implicit surface reconstruction, building upon the recent success of kernel methods for 3D reconstruction of oriented point clouds.
-As we show, both, from a theoretical and practical perspective, Matérn kernels have some appealing properties which make them particularly well suited for surface reconstruction—outperforming state-of-the-art methods based on the arc-cosine kernel while being significantly easier to implement, faster to compute, and scaleable.
-Being stationary, we demonstrate that the Matérn kernels’ spectrum can be tuned in the same fashion as Fourier feature mappings help coordinate-based MLPs to overcome spectral bias.
-Moreover, we theoretically analyze Matérn kernel’s connection to Siren networks as well as its relation to previously employed arc-cosine kernels.
+We propose to use the family of Matérn kernels for implicit surface reconstruction, building upon the recent success of kernel methods for 3D reconstruction of oriented point clouds.
+As we show from a theoretical and practical perspective, Matérn kernels have some appealing properties which make them particularly well suited for surface reconstruction—outperforming state-of-the-art methods based on the arc-cosine kernel while being significantly easier to implement, faster to compute, and scalable.
+Being stationary, we demonstrate that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias.
+Moreover, we theoretically analyze Matérn kernels’ connection to Siren networks as well as their relation to previously employed arc-cosine kernels.
 Finally, based on recently introduced Neural Kernel Fields, we present data-dependent Matérn kernels and conclude that especially the Laplace kernel (being part of the Matérn family) is extremely competitive, performing almost on par with state-of-the-art methods in the noise-free case while having a more than five times shorter training time.
@@
 Formally, let $\Omega\subset\mathbb{R}^{d}$ and $\mathcal{X}=\{x_{1},x_{2},\dots,x_{m}\}\subset\Omega$ be a set of $m$ points in $d$ dimensions which forms, together with associated per-point normals, a dataset $\mathcal{D}=\mathcal{X}\times\{n_{1},n_{2},\dots,n_{m}\}\subset\Omega\times\mathbb{R}^{d}$.
-The goal of surface reconstruction is to recover the objects’ shape from $\mathcal{D}$.
+The goal of surface reconstruction is to recover the object’s shape from $\mathcal{D}$.
 The shape of objects may be represented as dense point clouds, polygonal meshes, manifold atlases, voxel grids, or (as the zero-level set of) implicit functions, which is the representation we chose to focus on in this work.
@@
 $\hat{f}=\operatorname*{arg\,min}_{f\in H}\left\{\sum_{i=1}^{m}|f(x_{i})|^{2}+\|\nabla f(x_{i})-n_{i}\|^{2}+\lambda\|f\|_{H}^{2}\right\},$ (1)
-where the function space $H:=H(\Omega)$ is usually taken to be a reproducing kernel Hilbert space (RKHS) with associated reproducing kernel $k:\Omega^{2}\longrightarrow\mathbb{R}$, and $\lambda>0$.
+where the function space $H:=H(\Omega)$ is usually taken to be a Reproducing Kernel Hilbert Space (RKHS) with associated reproducing kernel $k:\Omega\times\Omega\longrightarrow\mathbb{R}$, and $\lambda>0$.
 Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (Lorensen & Cline, 1987).
 It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq.
-(1) turns into the following constrained minimization problem
+(1) turns into the constrained minimization problem,
 $\operatorname*{minimize}_{f\in H}\|f\|_{H}\quad\text{s.t.}\quad f(x_{i})=0\quad\text{and}\quad\nabla f(x_{i})=n_{i}\quad\forall i\in\{1,2,\dots,m\}$ (2)
-which makes it easier to see how $\hat{f}$ (thus the estimated surface $\mathcal{\hat{S}}$), behaves away from the input points, $\mathcal{X}$.
-Clearly, while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing.
-The properties of $\|f\|_{H}$ in $H$ are uniquely determined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq.
+which makes it easier to see how the predicted surface behaves away from the input points.
+As seen from Eq.
+(2), while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing.
+The properties of $\|f\|_{H}$ are uniquely defined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq.
 (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$.
 Here, $\alpha\in\mathbb{R}^{m}$ and $K=(K)_{ij}=k(x_{i},x_{j})\in\mathbb{R}^{m\times m}$.
-This shows that the behavior of the norm in Eq.
-(2) and hence, of estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by choosing different kernel functions, effectively allowing us to inject certain inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
+This shows that the behavior of the norm and hence, estimated surfaces $\mathcal{\hat{S}}$, can be explicitly controlled by the kernel function, enabling easy injection of inductive biases (such as smoothness assumptions) into the surface reconstruction problem.
 ((a))
@@
 (2018)) of the network.
-Figure 2: Tunable Matérn kernels lead to better surface reconstructions compared to the previously employed arc-cosine kernel.
-Here, we show surface reconstructions from sparse point clouds of just 1,000 points.
 One important aspect of this kernel is that it does not have adjustable parameters.
 While this can be advantageous in some situations, it also prevents users from tuning their reconstructions.
-Clearly, if the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
-In this work, we propose to use a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2.
-Contrary to arc-cosine kernels, Matérn kernels are stationary (hence rotation- and translation-invariant), locally-supported (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace (for $\nu=1/2$) and Gaussian kernel (as $\nu\rightarrow\infty$).
-As we will show, both, from a theoretical and practical perspective, Matérn kernels have some intriguing properties which make them particularly well suited for surface reconstruction.
-Most notably, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details).
-In addition, our theoretical analysis reveals that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activation function; Sitzmann et al.
+If the arc-cosine kernel fails to accurately recover a surface, then there is no possibility (except for using more observations) to improve the result.
+Figure 2: Tunable Matérn kernels lead to better surface reconstructions than the previously employed arc-cosine kernel.
+Here, we show surface reconstructions from sparse point clouds of just 1,000 points.
+In this work, we suggest a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2.
+Contrary to arc-cosine kernels, Matérn kernels are stationary (hence translation invariant), spatially decaying (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace and Gaussian kernel.
+As we will show, Matérn kernels have appealing properties, making them the ideal candidate for kernel-based surface reconstruction.
+From a practical perspective, we demonstrate that a simple change of the kernel function from arc-cosine to Matérn leads to a consistently improved reconstruction accuracy and a significant speedup.
+From a theoretical perspective, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details).
+In addition, we prove that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activations; Sitzmann et al.
 (2020)) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs.
-Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (i.e., the Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the hypersphere $\mathbb{S}^{d-1}$.
-Our experimental evaluation demonstrates that Matérn 1/2 and 3/2 are extremely competitive, outperforming the arc-cosine kernel while being significantly easier to implement (essentially two lines of standard PyTorch code), faster to compute, and scalable.
-Moreover, by leveraging Neural Kernel Fields (Williams et al., 2022), we show that data-dependent (a.k.a.
-learnable) Matérn kernels perform almost on par with highly sophisticated and well-engineered NKSR in the noise-free case while being more than five times faster to train.
+Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the sphere.
+In summary, our key contributions are:
+  • We propose to use Matérn kernels for tunable implicit surface reconstruction.
+  • We theoretically analyze Matérn kernels, relating them to Fourier feature mappings, Siren networks, and arc-cosine kernels.
+Moreover, we derive practical insights into how to choose the tunable bandwidth parameter based on a new bound of the $L_{2}$ reconstruction error.
+  • We propose data-dependent (i.e., learnable) Matérn kernels by leveraging the Neural Kernel Field (NKF) framework (Williams et al., 2022).
+Our experimental evaluation reveals that Matérn 1/2 and 3/2 are extremely competitive, outperforming the arc-cosine kernel while being significantly easier to implement (essentially two lines of standard PyTorch code), faster to compute, and scalable.
+In addition to geometry, we show that Matérn kernels surpass the arc-cosine kernel in reconstructing other high-frequency scene attributes, such as texture.
+Finally, we demonstrate that learnable Matérn kernels (1) outperform the data-dependent arc-cosine kernel (as implemented in the original NKF framework) while being more than four times faster to train, and (2) perform almost on par with highly sophisticated and well-engineered NKSR in the noise-free case while having a more than five times shorter training time.
 2 Related Work
@@
 (2022b).
-Early approaches for kernel-based 3D surface reconstruction mostly employ radial basis functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005).
-Nowadays, the probably most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; (Kazhdan & Hoppe, 2013)) which, however, can itself be viewed as kernel method as shown in Williams et al.
-(2021).
+Early kernel-based 3D surface reconstruction methods mostly employ Radial Basis Functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005).
+Nowadays, the most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; Kazhdan & Hoppe (2013)) which, however, can itself be viewed as kernel method (Williams et al., 2021).
 Only recently, Neural Splines (NS; Williams et al.
@@
 $k_{\text{AC}}(x,y)=\frac{\|x\|\|y\|}{\pi}\left(\sin\theta+(\pi-\theta)\cos\theta\right),\quad\text{where}\quad\theta=\cos^{-1}\left(\frac{x^{\top}y}{\|x\|\|y\|}\right)$ (3)
-for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques as well as non-linear methods based on neural networks by a large margin.
-This method laid the cornerstone for the Neural Kernel Field (NKF; Williams et al.
-(2022)) framework which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016).
-Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn a set of spatially-varying basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense.
+for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques and non-linear methods based on neural networks by a large margin.
+This method laid the cornerstone for Neural Kernel Fields (NKFs; Williams et al.
+(2022)) which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016).
+Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense.
 Neural Kernel Surface Reconstruction (NKSR; Huang et al.
-(2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose surface reconstruction method that is highly scaleable and robust against noise, eventually becoming state-of-the-art.
-3 Background: Matérn Kernels
-We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) instead of the arc-cosine kernel for implicit surface reconstruction, being defined as:
+(2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose, highly scalable surface reconstruction method that is robust against noise and eventually became state-of-the-art.
+3 Matérn Kernels for Tunable Surface Reconstruction
+We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) for implicit surface reconstruction, being defined as
 $k_{\nu}(x,y)=\Phi_{\nu}(\|x-y\|)=\Phi_{\nu}(\tau)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\tau}{h}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\tau}{h}\right),$ (4)
@@
 $\Gamma$ denotes the Gamma function, and $K_{\nu}$ is the modified Bessel function of the second kind of order $\nu$.
-For $\nu=p+\frac{1}{2}$, $p\in\mathbb{N}^{+}$ (i.e., if $\nu$ is a half-integer), Eq.
-(4) simplifies to a product of an exponential and polynomial of degree $p$, that is
-$\Phi_{p+\frac{1}{2}}(\tau)=\exp\left(-\frac{\sqrt{2p+1}\tau}{h}\right)\frac{p!}{(2p)!}\sum_{i=0}^{p}\frac{(p+i)!}{i!(p-i)!}\left(\frac{2\sqrt{2p+1}\tau}{h}\right)^{p-i}.$ (5)
-Matérn kernels generalize a variety of other kernels; using Eq. (5), the most popular ones can be written in closed-form as:
-$p$ $=0:\Phi_{\text{1/2}}(\tau)=\exp\left(-\frac{\tau}{h}\right),$ (6)
-$p$ $=1:\Phi_{\text{3/2}}(\tau)=\exp\left(-\frac{\sqrt{3}\tau}{h}\right)\left(1+\frac{\sqrt{3}\tau}{h}\right),$ (7)
-$p$ $=2:\Phi_{\text{5/2}}(\tau)=\exp\left(-\frac{\sqrt{5}\tau}{h}\right)\left(1+\frac{\sqrt{5}\tau}{h}+\frac{5\tau^{2}}{3h^{2}}\right),$ (8)
+Matérn kernels generalize a variety of other kernel functions; the most popular ones can be written in closed form as
+$\nu$ $=1/2:\Phi_{\text{1/2}}(\tau)=\exp\left(-\frac{\tau}{h}\right),$ (5)
+$\nu$ $=3/2:\Phi_{\text{3/2}}(\tau)=\exp\left(-\frac{\sqrt{3}\tau}{h}\right)\left(1+\frac{\sqrt{3}\tau}{h}\right),$ (6)
+$\nu$ $=5/2:\Phi_{\text{5/2}}(\tau)=\exp\left(-\frac{\sqrt{5}\tau}{h}\right)\left(1+\frac{\sqrt{5}\tau}{h}+\frac{5\tau^{2}}{3h^{2}}\right),$ (7)
 where $\Phi_{\text{1/2}}$ is known as the Laplace kernel. In the limiting case of $\nu\rightarrow\infty$, Matérn kernels recover the popular Gaussian kernel:
 $\Phi_{\infty}(\tau):=\lim_{\nu\rightarrow\infty}\Phi_{\nu}(\tau)=\exp\left(-\frac{\tau^{2}}{2h^{2}}\right).$ (8)
-The spectral density of Matérn kernels is given as the inverse Fourier transform of $\Phi_{\nu}$, that is
-$p_{\nu}(\omega)=\mathcal{F}^{-1}[\Phi_{\nu}]=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$ (10)
-which is due to the fact that Matérn kernels are stationary (or shift-invariant—they only depend on the difference $x-y$ between two points, $x,y\in\Omega$); hence, kernel and associated spectral density are Fourier duals (a consequence of Bochner’s Theorem).
-Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$.
-The RKHS induced by Matérn kernels is quite well understood (see, e.g., (Kanagawa et al., 2018)) and can be formulated as $H_{\nu}=\{f\in L_{2}(\mathbb{R}^{d})\cap C(\mathbb{R}^{d}):\|f\|_{H_{\nu}}^{2}<\infty\}$, where the RKHS norm is
-$\|f\|_{H_{\nu}}^{2}=h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}+\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega.$ (11)
-For more information about Matérn kernels, we refer the interested reader to Porcu et al. (2023) for a recent review. Next, we revisit some basic properties of Matérn kernels.
+For more information about Matérn kernels, please refer to Porcu et al. (2023). Next, we revisit some basic properties of Matérn kernels and place them in the context of surface reconstruction.
 3.1 Basic Properties
 Differentiable.
-Matérn kernels allow for controlled smoothness, being exactly $\lceil\nu\rceil-1$ times differentiable (in the mean-square sense); for instance, while the Matérn 1/2 kernel is zero times differentiable, the Gaussian kernel (for $\nu\rightarrow\infty$) is infinitely often differentiable.
+Matérn kernels allow for controlled smoothness, being exactly $\lceil\nu\rceil-1$ times differentiable (in the mean-square sense).
+Since functions $f$ in an RKHS $H$ inherit the differentiability class of the inducing kernel $k$ due to the reproducing property, the from Eq.
+(1) reconstructed surface $\mathcal{\hat{S}}$ enjoys the same smoothness properties as $k$.
+In the context of 3D surface reconstruction, this allows an easy injection of inductive biases into the reconstruction problem; for instance, if the roughness or noisiness of the objects to be reconstructed is known in advance, one can adjust the smoothness of the kernel accordingly.
+This is not possible with the arc-cosine kernel.
 Stationary.
 Matérn kernels are stationary, i.e., they only depend on the difference $x-y$ between two points $x,y\in\Omega$.
-If the distance is Euclidean, Matérn kernels are also isotropic (or homogeneous); they only depend on $\|x-y\|$.
-As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of $\mathcal{X}$.
-We note that, while the previously used arc-cosine kernel is also rotation invariant, it is not translation invariant.
-Locally supported.
-As opposed to arc-cosine kernels, Matérn kernels are locally supported (or compact); that is, they tend to zero as the distance $\|x-y\|$ becomes large.
-This is a very desirable property, especially from a computational perspective, as it leads to sparse Gram matrices when multiplied with the kernel $\Phi^{\prime}_{\nu^{\prime}}(\tau)=\max\left\{0,\left(1-\tau/h^{\prime}\right)^{\nu^{\prime}}\right\}$ for suitable $\nu^{\prime}$ and $h^{\prime}$, effectively allowing the use of highly efficient and scalable sparse linear solvers.
-Please see Genton (2001) on how to choose suitable values for $\nu^{\prime}$ and the cut-off distance, $h^{\prime}$.
+If the distance is Euclidean, Matérn kernels are also isotropic; they only depend on $\|x-y\|$.
+As a result, Matérn kernels are rotation and translation invariant, hence being independent of the absolute embedding of points $\mathcal{X}$ (based on the reproducing property, translating input points does not change the reconstructed surface).
+Consequently, since the kernel’s value only depends on the relative positions, objects in a scene (or parts of an object) with similar geometric properties will be reconstructed consistently as long as the relative distances between points on each instance (or part) remain the same.
+This is in contrast to the arc-cosine kernel, whose value depends on the absolute position of points, see Eq.
+(3); hence, it is not translation invariant (thus non-stationary), and multiple instances of an object in a scene will generally not yield the same surface reconstruction.
+See Appendix A for further discussion.
+Spatially decaying.
+As opposed to arc-cosine kernels, Matérn kernels are spatially decaying; they tend to zero as the distance $\|x-y\|$ becomes large.
+Consequently, although Matérn kernels are technically not compact (or locally supported), because their value decays exponentially fast (Porcu et al., 2023), they can be considered effectively compact.
+From a computational perspective, this is a very desirable property as it leads to sparse Gram matrices when truncated with a special kernel (Genton, 2001), allowing the use of highly efficient and scalable sparse linear solvers.
+Finally, we note that simply truncating any kernel does not yield a valid, positive definite kernel in general.
+We now derive new theoretical insights into Matérn kernels, ultimately aiming to provide arguments as to why we believe this family of kernels is particularly well suited for surface reconstruction.
 ((a))
@@
 While we fix $h=1$ and vary $\nu$ on the left, the EDR for $\nu=1/2$ and different values of $h$ is shown on the right.
-Larger values of $\nu$ and $h$, i.e., smoother kernels, lead to faster eigenvalue decay; hence slow convergence to high-frequency details.
-4 Matérn Kernels for Tunable Surface Reconstruction
-We now derive new theoretical insights into Matérn kernels, ultimately aiming to provide arguments as to why we believe this family of kernels is particularly well suited for surface reconstruction.
+Larger values of $\nu$ and $h$, i.e., smoother kernels, lead to faster eigenvalue decay; hence, slow convergence to high-frequency details.
 3.2 Eigenvalue Decay, Spectral Bias, and Reconstruction Error
 Eigenvalue decay and spectral bias.
-We begin by analyzing Matérn kernel’s eigenvalue decay rate (EDR) and show that it can be tuned by varying the smoothness parameter $\nu$ and/or bandwidth $h$.
+We begin by showing that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias.
 As investigated in Tancik et al.
-(2020) for coordinate-based MLPs, a rapid decrease of the associated NTK’s eigenvalues implies slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components).
-Consequently, a smaller EDR, i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content; hence, in the context of implicit surface reconstruction, more detailed geometry.
+(2020), a rapid decrease of the NTK’s eigenvalues implies an associated MLP’s slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components).
+Consequently, a smaller eigenvalue decay rate (EDR), i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content—in the context of implicit surface reconstruction, more detailed geometry.
 To overcome this slow convergence, a phenomenon known as spectral bias, Tancik et al.
-(2020) propose to use a (random) Fourier feature mapping of the form $\gamma(x)=[\cos(2\pi Bx),\sin(2\pi Bx)]$ applied to the inputs $x\in\mathbb{R}^{d}$ before passing them to the MLP, effectively transforming the MLP’s NTK, $k_{\text{NTK}}$, into a stationary kernel with tunable bandwidth parameter, $k_{\text{NTK}}(\gamma(x)^{\top}\gamma(y))=k_{\text{NTK}}(k(x-y))=:k^{\prime}_{\text{NTK}}(x-y)$.
-When $B\in\mathbb{R}^{q\times d}$ is drawn from a standard normal distribution (i.e., zero mean and unit variance), $k$ is an approximation of the Gaussian kernel (Rahimi & Recht, 2007).
-This allows explicit manipulation of the kernels’ spectrum, overcoming spectral bias.
+(2020) use a Fourier feature mapping of the form
+$\gamma(x)=[a_{1}\cos(2\pi b_{1}^{\top}x),a_{1}\sin(2\pi b_{1}^{\top}x),\dots,a_{q}\cos(2\pi b_{q}^{\top}x),a_{q}\sin(2\pi b_{q}^{\top}x)]^{\top}$ (9)
+applied to the inputs $x\in\mathbb{R}^{d}$ before passing them to the MLP, effectively transforming the MLP’s NTK, $k_{\text{NTK}}$, into a stationary kernel, $k_{\text{NTK}}(\gamma(x)^{\top}\gamma(y))=k_{\text{NTK}}(k_{\gamma}(x-y))=:k^{\prime}_{\text{NTK}}(x-y)$ whose spectrum can be tuned through manipulation of the Fourier basis frequencies, $b_{i}\in\mathbb{R}^{d}$, and corresponding coefficients, $a_{i}\in\mathbb{R}$, of the kernel $k_{\gamma}(x-y)=\sum_{i=1}^{q}a^{2}_{i}\cos(2\pi b^{\top}_{i}(x-y))$.
+This shows that $\gamma$ enables explicit control over the NTK’s EDR, overcoming spectral bias.
 Matérn kernels, being stationary, allow for the same degree of control over the EDR as a Fourier feature mapping; their spectrum can also be tuned (by varying $\nu$ and/or $h$, see Figure 1).
@@
 (2024) for general input domains and distributions.
-As seen, the smoother the kernel (i.e., the higher $\nu$) and the higher $h$, the faster the eigenvalues decay.
-This implies less detailed surface reconstruction (with a higher error) for smoother kernels and greater $h$; indeed, that is exactly what we observe in practice, see Figure 4.
+As seen, the smoother the kernel (i.e., the larger $\nu$) and the greater $h$, the faster the eigenvalues decay.
+This implies less detailed surface reconstruction (with a higher error) for smoother kernels and larger $h$; indeed, this is exactly what we observe in practice, see Figure 4.
 Contrary, by lowering both parameters, one can achieve faster convergence to high-frequency geometric details, effectively leading to more nuanced surface reconstruction and a smaller reconstruction error.
-Interestingly, the first-order arc-cosine kernel and Matérn 1/2 have the same EDR if $h=1$.
 Figure 4: Matérn kernels are tunable.
 Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3).
-However, setting $h$ too small ($<0.5$) results in over-fitting, while setting $h$ too big ($>10$) over-smooths (i.e., under-fits) the true surface.
+However, setting $h$ too small ($<0.5$) results in overfitting, while setting $h$ too big ($>10$) oversmooths (i.e., underfits) the true surface.
 This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
-Reconstruction error. We proceed by investigating the bandwidth’s influence on the reconstruction error, measured by the mean-squared error which is defined (and can be bounded) as:
+Reconstruction error. We proceed by investigating the bandwidth’s influence on the $L_{2}$ reconstruction error, which is defined (and can be bounded) as
 $\|f-\hat{f}\|_{L_{2}}=\left(\int_{\Omega}\left(f(x)-\hat{f}(x)\right)^{2}\mathrm{d}x\right)^{1/2}\leq C_{\mathcal{X},\Omega}^{(\nu+d)/2}\|f\|_{H_{\nu}},$ (10)
-where $C_{\mathcal{X},\Omega}$ is a constant that only depends on $\mathcal{X}$ and $\Omega$ and for $\nu<\infty$.
+where $C_{\mathcal{X},\Omega}$ is a constant that only depends on $\mathcal{X}$ and $\Omega$, and
+$\|f\|_{H_{\nu}}^{2}=h^{2\nu}\left((2\pi)^{d/2}C_{d,\nu}\right)^{-1}\int_{\mathbb{R}^{d}}\left(\frac{2\nu}{h^{2}}+\left(2\pi\|\omega\|\right)^{2}\right)^{\nu+d/2}|\mathcal{F}[f](\omega)|^{2}\mathrm{d}\omega$ (11)
+is the norm of the RKHS induced by a Matérn kernel with bandwidth $h$ and smoothness $\nu$.
+Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$.
 For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., Santin & Schaback (2016).
@@
 where $h>0$, and $C^{1}_{d,\nu}$ and $C^{2}_{d,\nu}$ are functions of $\mathcal{F}[f]$ that do not depend on $h$.
-A proof can be found in Appendix A.
+A proof can be found in Appendix B.
 We immediately see that $\|f\|^{2}_{H_{\nu}}\rightarrow\infty$ in either cases, $h\rightarrow 0$ and $h\rightarrow\infty$.
@@
 $h^{*}=\left(\frac{d}{2\nu}Q\right)^{1/(2\nu+d)}\quad\text{with}\quad Q=\frac{C^{1}_{d,\nu}(\mathcal{F}[f])}{C^{2}_{d,\nu}(\mathcal{F}[f])}$ (12)
-(see Appendix B for a derivation). This can also be seen from the inset figure on the right, where we plot the RKHS norm with respect to the obtained bound as a function of $h$.
-Interestingly, our analysis reveals that the reconstruction error can not be made arbitrarily small by just lowering $h$; if the bandwidth $h$ is chosen too small, the reconstruction error starts to increase again.
-Conversely, if $h$ is set too high, the resulting surface is too smooth, ultimately leading to increased errors.
-In fact, what we actually observe is that Matérn kernels under-fit the true surface when $h$ is set too big, and, conversely, over-fit if $h$ is chosen too small, see also Figure 4.
-As we have shown, the optimal trade-off is reached if $h$ equals $h^{*}$.
-Moreover, whilst the bandwidth $h$ does play a critical role, our results show it is not very sensitive in practice, especially for Matérn 1/2 and 3/2 (see inset figure).
+Figure 5: RKHS norm as a function of $h$. We plot the bound from Proposition 2 as a function of $h$.
+(see Appendix C for details).
+Our analysis shows that the reconstruction error can not be made arbitrarily small by just lowering $h$; if $h$ is chosen too small, the reconstruction error starts to increase again, i.e., it overfits the true surface.
+Conversely, if $h$ is set too high, the resulting surface is too smooth (underfits the true surface), ultimately leading to increased errors.
+The optimal trade-off is reached if $h=h^{*}$.
+This is also observed in practice, see Figure 4.
+As seen from Figure 5, however, the described effect is more pronounced for smoother Matérn kernels (large $\nu$), as, at some point, the norm (hence, the reconstruction error) increases rapidly for large $h$.
+Our theoretical analysis reveals several practical insights.
+First, the reconstruction error is generally more sensitive to very small values of $h$.
+Second, a good starting point is always $h=1$.
 3.3 Relation to Neural Networks and the Arc-cosine Kernel
@@
 Next, we study the connection between Matérn kernels and neural networks.
-It is well known that first-order arc-cosine kernels mimic the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al.
+It is well known that the first-order arc-cosine kernel mimics the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al.
 (2021).
 Here, $[x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
-Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens Sitzmann et al.
-(2020) (fully-connected networks with sine activation function) if the layer width approaches infinity.
-We claim the following:
+Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens (Sitzmann et al., 2020) if the layer-width approaches infinity. We claim the following:
 Theorem 3.
 Consider a two-layer fully-connected network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ with sine activation function, $m$ hidden neurons, and fixed bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$.
-Let $h,\nu>0$ be given.
+Let $h,\nu>0$ and $C_{d,\nu}$ be defined as in Eq.
+(11).
 If $w_{i}$ is randomly initialized from
@@
 and $b_{i}\sim\mathcal{U}(0,2\pi)$, the NTK of $f$ is a Matérn kernel with bandwidth $h$ and smoothness $\nu$ as $m\rightarrow\infty$.
-A proof can be found in Appendix C. This result establishes for the first time a connection between Matérn kernels and the computation in neural networks, in particular, Sirens.
-Matérn kernels vs. arc-cosine kernels. Lastly, we aim to compare Matérn kernels with the previously used first-order arc-cosine kernels. Our analysis is based on:
+A proof can be found in Appendix D.
+This result establishes for the first time a connection between Sirens and kernel methods.
+While explicitly shown for Matérn kernels, as we argue in the Appendix D, the arguments of Theorem 3 in fact apply to all stationary kernels, making it a powerful tool to study the connection between widely used Sirens and any stationary kernel function.
+Matérn kernels vs. arc-cosine kernels. Lastly, we aim to compare Matérn kernels with the previously used first-order arc-cosine kernel. Our analysis is based on:
 Theorem 4 (Chen & Xu (2021), Theorem 1).
@@
 In other words, Theorem 4 shows that the RKHS of the Laplace kernel (a Matérn kernel with smoothness $\nu=1/2$) is the same as the RKHS induced by the NTK of a fully-connected ReLU network when inputs are restricted to $\mathbb{S}^{d-1}$.
-This immediately raises the question of whether the NTK for general input domains is significantly different from a standard Laplace kernel.
-Indeed, empirical results (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020) already suggest a negative answer; however, from a theoretical perspective, the claim is still unproven.
 Based on Theorem 4, we obtain the following connection between Laplace and the first-order arc-cosine kernel:
@@
 When restricted to the hypersphere $\mathbb{S}^{d-1}$, (1) the RKHS of the Matérn kernel with smoothness $\nu=1/2$ (the Laplace kernel) is equal to the RKHS induced by the first-order arc-cosine kernel, implying that (2) the Laplace and arc-cosine kernel are equal up to an affine transformation.
-See Appendix D for a proof.
-For general input domains, our empirical results presented next support the claim mentioned above (at least for ($L=2$)-layer networks), showing that the Laplace kernel even outperforms the arc-cosine kernel.
+See Appendix E for a proof.
+For general input domains, our empirical results presented next support (at least for ($L=2$)-layer networks) the widely believed claim that the NTK is not significantly different from standard kernels (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020), such as the Laplace kernel.
+Indeed, we show that the Laplace kernel even outperforms the arc-cosine kernel.
 4 Experiments and Results
@@
 (2013)) in Section 4.1.
-For ShapeNet, we use the train/val/test split provided by Mescheder et al.
-(2019).
+Moreover, Section 4.2 demonstrates Matérn kernels’ ability to reconstruct high-frequency textures on the Google Scanned Objects (GSO; Downs et al.
+(2022)) dataset and Objaverse (Deitke et al., 2023).
 Leveraging Neural Kernel Fields (Williams et al., 2022), we present and evaluate learnable Matérn kernels in Section 4.3.
@@
 ShapeNet.
-We compare Matérn kernels in a sparse setting against the arc-cosine kernel as implemented in Neural Splines (NS, Williams et al.
+We compare Matérn kernels in a sparse setting against the arc-cosine kernel on ShapeNet (using train/val/test split provided by Mescheder et al.
 (2019)).
@@
 We implemented Matérn kernels in PyTorch and took the official CUDA implementation of the arc-cosine kernel from NS, eventually integrated into a unified framework to ensure fair comparison.
-Notably, we did not use Nyström sampling as in NS, and employ PyTorch’s direct Cholesky solver to solve the KRR problem in Eq.
+Notably, we did not use Nyström sampling as in NS and employed PyTorch’s direct Cholesky solver to solve the KRR problem in Eq.
 (1) instead of an iterative conjugate gradient solver.
@@
 (1) with finite differences.
-See Appendix E for details.
-F-Score $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 Matérn 1/2 93.6 93.7 93.7 4.05 4.02 4.02 Matérn 3/2 94.6 94.8 94.9 4.05 4.00 4.09 Matérn 5/2 92.4 93.8 92.9 6.42 5.65 6.91 Matérn $\infty$ 59.3 52.4 40.9 57.56 50.59 48.38 Arc-cosine 92.8 4.67 NS 90.6 4.74 SPSR 84.3 6.26       CD $\downarrow$ HD $\downarrow$ Time Matérn 1/2 0.21 4.43 9.85 sec Matérn 3/2 0.18 2.93 12.47 sec Matérn 5/2 0.58 22.52 15.26 sec Matérn $\infty$ 3.83 33.27 12.76 sec NS* 0.19 3.19 19.89 sec NS 0.17 2.85 11.91 sec SPSR 0.21 4.69 1.65 sec FFN 0.28 4.45 337.91 sec Siren 0.19 3.86 599.00 sec
+See Appendix F for details.
+F-Score $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 Matérn 1/2 93.6 93.7 93.7 4.05 4.02 4.02 Matérn 3/2 94.6 94.8 94.9 4.05 4.00 4.09 Matérn 5/2 92.4 93.8 92.9 6.42 5.65 6.91 Matérn $\infty$ 59.3 52.4 40.9 57.56 50.59 48.38 Arc-cosine 92.8 4.67 NS 90.6 4.74 SPSR 84.3 6.26      CD $\downarrow$ HD $\downarrow$ Matérn 1/2 0.21 4.43 Matérn 3/2 0.18 2.93 Matérn 5/2 0.58 22.52 Matérn $\infty$ 3.83 33.27 NS* 0.19 3.19 NS 0.17 2.85 SPSR 0.21 4.69 FFN 0.28 4.45 Siren 0.19 3.86 SAP 0.21 4.85 DiGS 0.18 3.55 VisCo 0.18 2.95 OG-INR 0.20 4.06
 […]
@@
 Matérn 5/2 and the Gaussian kernel (for $\nu\rightarrow\infty$) perform significantly worse, being simply too smooth.
-Figure 2 and Appendix E shows qualitative results.
+Figure 2 and Appendix F shows qualitative results.
+Time
+Matérn 1/2 9.85
+Matérn 3/2 12.47
+Matérn 5/2 15.26
+Matérn $\infty$ 12.76
+NS* 19.89
+NS 11.91
+SPSR 1.65
+Table 2: Runtime comparison on SRB. Time in seconds.
 SRB.
-Additionally, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark (Berger et al., 2013) which consists of five complex shapes simulated from noisy range scans.
-For this, we integrated Matérn kernels into the NS framework, using KeOps (Charlier et al., 2021) and the FALKON (Rudi et al., 2017) solver.
-We used the same parameter settings as in the original NS paper.
-In addition, for Matérn kernels, we did a modest parameter sweep over the bandwidth, $h\in\{0.5,1,2\}$, and took the reconstruction with the lowest Chamfer distance.
-Figure 5: Qualitative results for non-learnable kernels on SRB (left) and learned kernels on ShapeNet (right). Please see Appendix F and G for more qualitative results.
-Quantitative results obtained by employing Chamfer and Hausdorff distance (HD) can be found in Table 1, demonstrating that Matérn 3/2 performs on par with NS while being faster to compute (compared to what we measured for NS).
-See Figure 5 (left) and Appendix F for qualitative results and per-object metrics.
+Next, we evaluate Matérn kernels on the challenging Surface Reconstruction Benchmark which consists of five complex shapes simulated from incomplete and noisy range scans with up to 100,000 points.
+For this, we implemented a highly scalable version of Matérn kernels, closely following the NS framework to solve the KRR problem in Eq.
+(1) using Nyström sampling (with 15,000 samples) and the FALKON (Rudi et al., 2017) solver.
+We compare against NS and SPSR, as well as kernel-free methods, Fourier Feature Networks (FFNs; Tancik et al.
+(2020)), Siren, SAP (Peng et al., 2021), DiGS (Ben-Shabat et al., 2022), VisCo (Pumarola et al., 2022), and OG-INR (Koneputugodage et al., 2023).
+All methods are optimization-based (FFN and Siren are overfitted to a single shape) and use surface normals.
+For Matérn kernels, we did a modest parameter sweep over $h\in\{0.5,1,2\}$ and took the reconstruction with the lowest Chamfer distance.
+Runtime is measured on a single NVIDIA V100.
+Figure 6: Qualitative results for non-learnable kernels on SRB (left) and learned kernels on ShapeNet (right). Please see Appendices G and I for more qualitative results.
+Quantitative results obtained by employing Chamfer and Hausdorff distance (HD) can be found in Table 1, demonstrating that Matérn 3/2 performs on par with NS while being significantly faster to compute (compared to what we measured for NS), see Table 2.
+Moreover, Matérn 3/2 outperforms all evaluated state-of-the-art kernel-free methods (FFN, Siren, SAP, DiGS, VisCo, and OG-INR).
+Figure 6 (left) and Appendix G provide more qualitative results and per-object metrics.
+4.2 Texture Reconstruction on GSO and Objaverse
+$m$ Kernel PSNR $\uparrow$ LPIPS $\downarrow$ 10,000 Matérn 3/2 19.61 2.05 Arc-cosine 19.34 2.07 2,500 Matérn 3/2 18.92 2.15 Arc-cosine 18.88 2.09
+Table 3: Texture reconstruction on GSO. Matérn kernels outperform the arc-cosine kernel in dense and sparse settings.
+Lastly, we demonstrate Matérn kernels’ ability to represent other high-frequency scene attributes, such as texture.
+To do so, we randomly sample $m$ surface points with corresponding normals and per-point RGB color values from textured meshes, yielding an extended dataset $\mathcal{D}^{\prime}=\mathcal{D}\times\{c_{1},c_{2},\dots,c_{m}\}\subset\Omega\times\mathbb{R}^{d}\times\mathbb{R}^{3}$ for each shape.
+Then, instead of modeling an SDF $f$ as in Eq.
+(1), we are seeking a function $f^{\prime}:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{4}$ that, along with signed distances, also predicts per-point RGB values.
+Please find more information in Appendix H about how we adapt the KRR problem in Eq.
+(1) to estimate $\hat{f}^{\prime}$.
+Finally, we extract the object’s surface using Marching Cubes and trilinearily interpolate RGB values at previously predicted surface points.
+Figure 7: Texture reconstruction on Objaverse. Matérn kernels lead to fewer artifacts and sharper details than the arc-cosine kernel.
+We present quantitative results in terms of PSNR and LPIPS on GSO in Table 3 (metrics evaluated on the texture atlas; see Appendix H for further details).
+Matérn 3/2 surpasses the arc-cosine kernel in the densely sampled setting using 10,000 surface points, as well as in the sparse setting.
+Notably, we did not tune Matérn kernels’ bandwidth parameter for this experiment.
+We also show qualitative results on the challenging Objaverse (Deitke et al., 2023) dataset in Figure 7.
+We chose Objaverse as it includes extremely high-resolution and complex textures.
+Matérn 3/2 reconstructs high-frequency texture details with great precision, overall yielding visually more pleasant reconstructions than the arc-cosine kernel.
+Reconstructed textures have fewer artifacts and are generally sharper, demonstrating that Matérn kernels can overcome spectral bias much better than the arc-cosine kernel.
+Note that this perfectly confirms our theoretical analysis presented in Section 3.2.
 4.3 Data-dependent Matérn Kernels
@@
 We set $q=32$ for our experiments.
-Please refer to (Peng et al., 2020; Williams et al., 2022) for more details.
+Please see Williams et al.
+(2022) for details.
 Sparse surface reconstruction and extreme generalization.
-We compare learned Matérn kernels against the arc-cosine kernel originally used within the NKF framework and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet.
+We compare learned Matérn kernels against the original NKF framework (with the arc-cosine kernel) and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet.
 Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels.
-Additionally, we evaluate Matérn kernel’s out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories.
+Additionally, we evaluate Matérn kernels’ out-of-category generalization ability in an extreme setting, in which we train only on chairs and evaluate on the other 12 ShapeNet categories.
 Table 4 reports the results, quantified using intersection-over-union (IoU), Chamfer distance, and normal consistency (NC).
@@
 Regarding out-of-category generalization, we observe that Matérn 1/2 performs best, followed by Matérn 3/2, NKF, and NKSR.
-Please see Figure 5 (right) and Appendix G for qualitative results.
+Please see Figure 6 (right) and Appendix I for qualitative results.
 IoU $\uparrow$ CD $\downarrow$ NC $\uparrow$ Matérn 1/2 95.3 2.58 95.6 Matérn 3/2 94.9 2.70 95.3 Matérn 5/2 93.3 3.07 94.9 Matérn $\infty$ 92.1 3.39 94.2 NKF* 94.7 2.70 95.2 NKF 94.7 2.65 95.4 NKSR 95.6 2.34 95.4       Train only on chairs, test on all IoU $\uparrow$ CD $\downarrow$ NC $\uparrow$ Time/epoch Matérn 1/2 93.4 3.06 94.3 7.71 min Matérn 3/2 92.8 3.34 94.2 8.26 min Matérn 5/2 90.5 4.11 93.9 8.56 min Matérn $\infty$ 84.7 6.70 92.6 7.74 min NKF* 92.8 3.30 94.1 31.98 min NKSR 89.6 2.70 94.1 41.66 min
@@
 Table 5: Robustness against noise.
-We compare Matérn kernels’ robustness against noise on a subset of the ShapeNet dataset (including about 1,700 shapes) using different noise levels, $\sigma\in\{0,0.0025,0.005\}$ and bandwidths, $h\in\{0.5,1,2\}$.
-Again, NKF* denotes a re-implementation of NKF.
+We compare Matérn kernels’ robustness against noise on a subset of the ShapeNet dataset using different noise levels, $\sigma\in\{0,0.0025,0.005\}$ and bandwidths, $h\in\{0.5,1,2\}$.
 Bold marks best result, underline second best.
@@
 Convergence speed.
-Finally, we analyze Matérn kernel’s convergence speed in comparison to the arc-cosine kernel as implemented in NKF, see inset figure on the right.
+Finally, we analyze Matérn kernels’ convergence speed in comparison to the arc-cosine kernel as implemented in NKF, see inset figure on the right.
 As exemplified, Matérn 1/2 already converges after just 100 epochs, while the arc-cosine kernel requires twice as many epochs, also never being able to reach the same top accuracy.
@@
 5 Conclusion
-In this work, we have proposed to use the family of Matérn kernels for tunable implicit surface reconstruction, acting as a drop-in replacement of the recently employed first-order arc-cosine kernel within frameworks such as Neural Splines (Williams et al., 2021) and Neural Kernel Fields (Williams et al., 2022).
-We presented an in-depth theoretical analysis of Matérn kernels, analyzing its relation to Fourier feature mappings, Siren networks, as well as arc-cosine kernels.
-Our systematic evaluation confirmed our theoretical results and showed that Matérn kernels are significantly faster to compute and train, and, especially Matérn 1/2 and 3/2, lead to better surface reconstructions compared to the arc-cosine kernel—both, in a non-learnable as well as learnable regime.
-In the future, we plan to incorporate Matérn kernels into the NKSR framework to enhance robustness against noise.
+In this work, we have proposed to use the family of Matérn kernels for implicit surface reconstruction and showed that it consistently outperforms the recently employed first-order arc-cosine kernel—both, in a non-learnable as well as learnable regime—while being significantly easier to implement, faster to compute and train, and highly scalable.
+We demonstrated that Matérn kernels lead to tunable surface reconstruction, and, based on a new bound of the $L_{2}$ reconstruction error, derive practical insights into how to choose their tunable bandwidth parameter.
+Moreover, we presented an in-depth theoretical analysis, analyzing Matérn kernels’ relation to widely used Fourier feature mappings, Siren networks, and arc-cosine kernels.
 Acknowledgments
@@
 The authors are responsible for the content of this publication.
-Appendix A Proof of Proposition 2
+The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b112dc IRRW.
+NHR funding is provided by federal and Bavarian state authorities.
+NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683.
+Appendix A On the Stationarity of Matérn Kernels
+As noted in Section 3.1 of the main paper, Matérn kernels are stationary: they only depend on the difference $x-y$ between two points $x,y\in\Omega$, hence being translation invariant.
+This is because $k(x-y)=k(x+t-(y+t))$ for $t\in\mathbb{R}^{d}$ and all stationary kernels, $k$.
+In contrast, the arc-cosine kernel is not translation invariant as its value depends on the absolute positions, $x,y$, see Eq.
+(3) of the main paper (i.e., the arc-cosine kernel can not be written as a function of $x-y$).
+In fact, the arc-cosine kernel is non-stationary.
+In the following and with the help of Figure 8, we will further discuss why stationary kernels, such as Matérn kernels, might be a better choice for kernel-based implicit surface reconstruction than non-stationary kernels, such as the arc-cosine kernel.
+First, stationary kernels are naturally translation invariant.
+Consequently, they are independent of the absolute embedding of the input points—translating them does not change the reconstructed surface as the kernel’s value only depends on the relative distances between points.
+This is highly beneficial from a practical perspective since no additional steps (such as centering the point cloud) have to be taken before surface reconstruction.
+Being non-stationary, the arc-cosine kernel behaves differently.
+If we would just use the kernel as in Eq.
+(3) of the main paper, the reconstructed surface would vary depending on the actual location of the input points in space.
+This is a highly undesirable property when it comes to surface reconstruction.
+To make the arc-cosine kernel translation invariant, input points must be centered before reconstruction (thus eliminating the effects of translation).
+A centered arc-cosine kernel, denoted as $\bar{k}_{\text{AC}}$, reads:
+$\bar{k}_{\text{AC}}(x,y)=\frac{\|x-\bar{c}\|\|y-\bar{c}\|}{\pi}\left(\sin\bar{\theta}+(\pi-\bar{\theta})\cos\bar{\theta}\right),\quad\text{where}\quad\bar{\theta}=\cos^{-1}\left(\frac{(x-\bar{c})^{\top}(y-\bar{c})}{\|x-\bar{c}\|\|y-\bar{c}\|}\right).$
+Here, $c\in\mathbb{R}^{d}$ denotes the centroid of the input point cloud.
+$\bar{k}_{\text{AC}}$ is obviously translation invariant because $\|x+t-(\bar{c}+t)\|\|y+t-(\bar{c}+t)\|=\|x-\bar{c}\|\|y-\bar{c}\|$ and $(x+t-(\bar{c}+t))^{\top}(y+t-(\bar{c}+t))=(x-\bar{c})^{\top}(y-\bar{c})$ for all $t\in\mathbb{R}^{d}$.
+This follows because translating the input points also shifts its centroid.
+In conclusion, while stationary Matérn kernels can be directly used for kernel-based surface reconstruction, additional pre-processing steps are necessary for the arc-cosine kernel.
+Figure 8: Matérn kernels are translation invariant and lead to locally consistent surface reconstructions.
+Here, we show that two identical chairs in a scene are reconstructed differently depending on the absolute position of the input points when using the non-stationary arc-cosine kernel (especially noticeable at the chair’s legs).
+In contrast, stationary Matérn kernels lead to locally consistent surface reconstructions, being independent of the actual embedding (i.e., both chairs have the exact same reconstructed surface).
+Second, stationary kernels lead to locally consistent surface reconstructions.
+Since a stationary kernel’s value only depends on the relative positions, objects in a scene (or parts of objects) with similar geometric properties will be reconstructed consistently as long as the relative distances between points on each instance (or part) remain the same.
+This is a crucial and very desirable property for surface reconstruction which is not shared by the arc-cosine kernel (because it is non-stationary).
+Appendix B Proof of Proposition 2
 See 2
@@
 which concludes the proof. ∎
-Appendix B Derivation of $h^{*}$ in Eq. (13)
+Appendix C Derivation of $h^{*}$ in Eq. (12)
 Taking derivative of the bound presented in Proposition 2 w.r.t. $h$ yields
@@
 which shows what we have stated in Eq. (12) of the main paper.
-Appendix C Proof of Theorem 3
+Appendix D Proof of Theorem 3
 The proof of Theorem 3 is based on Bochner’s theorem (in harmonic analysis) which reads:
@@
 Choose
-$w_{i}\sim p_{\nu}(\omega)=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$
+$w_{i}\sim p_{\nu}(\omega)=h^{-2\nu}C_{d,\nu}\left(\frac{2\nu}{h^{2}}+(2\pi\|\omega\|)^{2}\right)^{-(\nu+d/2)}$ (13)
 for $h,\nu>0$ and $b_{i}\sim\mathcal{U}(0,2\pi)$ for all $i\in\{1,2,\dots,n\}$.
@@
 $k_{\text{NTK}}(x,y)$ $=\mathbb{E}_{(w,b)}\left[2\cos(w^{\top}x+b)\cos(w^{\top}y+b)\right]-2\underbrace{\mathbb{E}_{(w,b)}\left[\cos(w^{\top}(x+y)+2b)\right]}_{=0\,(*)}$
-${=}k_{\nu}(x,y)$
-where ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see Eq.
-(10) of the main paper.
+${=}k_{\nu}(x,y).$
+Here, ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see, e.g., Eq.
+(10) in Kanagawa et al.
+(2018).
 ∎
-Appendix D Proof of Corollary 5
+In fact, Theorem 3 can be generalized in that the NTK of a Siren is generally any stationary kernel:
+Remark 7.
+The NTK of a Siren can be associated with any stationary kernel if the bottom-layer weights in Eq. (13) are initialized according to the corresponding kernel’s spectral density.
+This is a direct consequence of Bochner’s theorem.
+Appendix E Proof of Corollary 5
 We first re-state the following result, needed to prove the second part of Corollary 5:
@@
 This shows that the Laplace and arc-cosine kernel are indeed related by an affine transformation when restricted to $\mathbb{S}^{d-1}$, concluding the proof. ∎
-Appendix E Additional Details for ShapeNet Experiment
-In this section, we provide additional details on the ShapeNet evaluation as presented in Section 5.1 of the main paper.
-E.1 Implementation Details
+Appendix F Additional Details for ShapeNet Experiment
+In this section, we provide implementation details and additional qualitative results on the ShapeNet experiment as presented in Section 4.1 of the main paper.
+F.1 Implementation Details
 Following common practice (see, e.g., Williams et al.
@@
 (1) is of the form
-$f(x)=\sum_{x^{\prime}\in\mathcal{X}^{\prime}}\alpha_{i}k(x,x^{\prime})$
+$f(x)=\sum_{i=1}^{2m}\alpha_{i}k(x,x_{i}^{\prime})$
 which is linear in the coefficients $\alpha\in\mathbb{R}^{2m}$, given by
@@
 and
-$y_{i}=\begin{cases}+{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{+},\\ -{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{-}.\end{cases}$
-$I\in\mathbb{R}^{2m\times 2m}$ denotes the identity matrix. We employ PyTorch’s built-in Cholesky solver to numerically stable solve for $\alpha$ and set ${\epsilon}=0.005$ for all experiments.
-E.2 Further Qualitative Results
+$y_{i}=\begin{cases}+{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{+},\\ -{\epsilon}&\text{if }x^{\prime}_{i}\in\mathcal{X}^{-},\end{cases}\quad\text{for }i=1,2,\dots,2m.$
+Here, $I\in\mathbb{R}^{2m\times 2m}$ denotes the identity matrix. We employ PyTorch’s built-in Cholesky solver to numerically stable solve for $\alpha$ and set ${\epsilon}=0.005$ for all experiments.
+F.2 Further Qualitative Results
 We show additional qualitative results on ShapeNet in Figure 9.
-Figure 6: More qualitative results on ShapeNet, comparing Matérn 1/2 and 3/2 for different values of $h$ to the arc-cosine kernel as implemented in Neural Splines (Williams et al., 2021).
-Appendix F Further Details on the Surface Reconstruction Benchmark
-F.1 Experimental Setting
-For Neural Splines (NS; Williams et al.
-(2021)), we used the official implementation provided here: https://github.com/fwilliams/neural-splines.
+Figure 9: Additional qualitative results on ShapeNet for non-learnable kernels. We compare the arc-cosine kernel against Matérn 1/2 and 3/2 for different values of $h$.
+Appendix G Further Details on the Surface Reconstruction Benchmark
+This section details the experimental setting that has been used to run Neural Splines (NS; Williams et al.
+(2021)) on SRB in Section 4.1 of the main paper.
+Moreover, we present per-object metrics and additional qualitative results.
+G.1 Experimental Setting
+For NS, we used the official implementation provided here: https://github.com/fwilliams/neural-splines.
 Similar to the original paper, we used 15,000 Nyström samples and did the same parameter sweep over the regularization parameter, $\lambda\in\{0,10^{-13},10^{-12},10^{-11},10^{-10}\}$.
@@
 Despite our best efforts, we were unable to reproduce the results for the NS kernel reported in the original paper.
-We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines) and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance.
-In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; the authors of (Williams et al., 2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively.
+We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines), and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance.
+In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; Williams et al.
+(2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively.
 In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
@@
 Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters.
-Runtime was measured on a single NVIDIA V100, similar to (Williams et al., 2021).
-F.2 Additional Results
-Per-object metrics are shown in Table 4, and more qualitative result in Figure 7.
-Figure 7: Qualitative results on the Surface Reconstruction Benchmark, comparing Matérn 1/2 and 3/2 to Neural Splines (NS; Williams et al. (2021)).
+Runtime was measured on a single NVIDIA V100, similar to Williams et al.
+(2021).
+G.2 Additional Results
+Per-object metrics are shown in Table 6, and more qualitative results in Figure 10.
 Anchor Daratech DC Gargoyle Lord Quas
@@
 Siren 0.32 8.19 0.21 4.30 0.15 2.18 0.17 4.64 0.17 0.82
-Table 4: Per-object quantitative results on the Surface Reconstruction Benchmark.
-Appendix G Additional Results for Data-dependent Kernels
+SAP 0.34 8.83 0.22 3.09 0.17 3.30 0.18 5.54 0.13 3.49
+DiGS 0.28 5.71 0.21 5.02 0.15 2.13 0.16 3.81 0.12 1.10
+VisCo 0.21 3.00 0.26 4.06 0.15 2.22 0.17 4.40 0.12 1.06
+OG-INR 0.29 7.56 0.23 2.89 0.17 2.68 0.19 5.01 0.13 2.14
+Table 6: Per-object quantitative results on SRB.
+”NS*” denotes the best result we could achieve by running the official implementation of NS, see Section G.1.
+Bold marks best result, underline second best.
+Figure 10: Additional qualitative results on SRB for non-learnable kernels. We compare Neural Splines (NS; Williams et al. (2021)) to Matérn 1/2 and 3/2.
+Appendix H Additional Details For Texture Reconstruction Experiment
+This section provides implementation details and further information about the evaluation protocol used for the texture reconstruction experiment presented in Section 4.2 of the main paper.
+H.1 Implementation Details
+Recall from Section 4.2 of the main paper that we are given a dataset $\mathcal{D}^{\prime}=\mathcal{D}\times\{c_{1},c_{2},\dots,c_{m}\}\subset\Omega\times\mathbb{R}^{d}\times\mathbb{R}^{3}$ with per-point RGB colors $\{c_{1},c_{2},\dots,c_{m}\}$.
+In order to also reconstruct texture, we are seeking a function $f^{\prime}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{4}$, predicting signed distances as well as RGB values.
+Using notation introduced in Section F.1, we compute $\alpha^{\prime}\in\mathbb{R}^{2m\times 4}$ as
+$\alpha^{\prime}=(K+\lambda I)^{-1}y^{\prime},\quad\text{where}\quad y^{\prime}:=[y,\bar{c}]\in\mathbb{R}^{2m\times 4}$
+and $\bar{c}:=[c,c]^{\top}$ with $(c_{1},c_{2},\dots,c_{m})\in\mathbb{R}^{m\times 3}$. Then, we predict the function $f^{\prime}$ as
+$f^{\prime}(x)=\sum_{i=1}^{2m}\alpha_{i}^{\prime}k(x,x_{i}^{\prime}),$
+where $\alpha_{i}^{\prime}\in\mathbb{R}^{4}$ denotes the $i$-th row of $\alpha^{\prime}$.
+Clearly, $f^{\prime}(x)=(f_{1}^{\prime}(x),f_{2}^{\prime}(x),f_{3}^{\prime}(x),f_{4}^{\prime}(x))\in\mathbb{R}^{4}$ with $f^{\prime}_{1}$ denoting the singed distance at $x$, and $f^{\prime}_{2},f^{\prime}_{3},f^{\prime}_{4}$ are the component functions that represent RGB colors at a position $x$.
+To extract a textured surface mesh from the predicted four-dimensional volume, we use a two-stage process.
+First, we apply Marching Cubes (Lorensen & Cline, 1987) to $f^{\prime}_{1}$, and then trilinearily interpolate RGB colors at the previously predicted surface points.
+H.2 Evaluation Protocol
+Our quantitative evaluation on the Google Scanned Objects (GSO; Downs et al.
+(2022)) dataset closely follows Mitchel et al.
+(2024), in which authors compute image-based evaluation metrics on the texture atlas to quantify reconstruction ability.
+Specifically, we first sample $m$ surface points along with corresponding normals and per-point RGB colors from shapes in GSO.
+Next, we solve the KRR problem in Eq.
+(1) as always.
+Then, instead of predicting RGB colors on the estimated geometry (which generally differs between various kernel functions), we predict (i.e., trilinearily interpolate) colors on the surface points of the original mesh.
+This allows us to leverage the existing UV parametrization of the original mesh to generate a texture atlas with colors predicted from KRR.
+Once the texture atlas at hand, we compare it to the original texture atlas using PSNR and LPIPS.
+We used ten objects from GSO with different complexity. The following assets have been used:
+  • ACE_Coffee_Mug_Kristen_16_oz_cup
+  • Animal_Planet_Foam_2Headed_Dragon
+  • Jansport_School_Backpack_Blue_Streak
+  • Now_Designs_Bowl_Akita_Black
+  • Predito_LZ_TRX_FG_W
+  • Reebok_ZIGTECH_SHARK_MAYHEM360
+  • Razer_Kraken_Pro_headset_Full_size_Black
+  • RJ_Rabbit_Easter_Basket_Blue
+  • Schleich_Lion_Action_Figure
+  • Weisshai_Great_White_Shark
+They can be downloaded from: https://app.gazebosim.org/dashboard.
+Appendix I Additional Results for Data-dependent Kernels
 We show additional qualitative results for learned kernels on ShapeNet in Figure 11.
-Figure 8: Additional qualitative results for learned kernels on ShapeNet.
-We compare against Neural Kernel Fields (NKFs; Williams et al.
+Figure 11: Additional qualitative results on ShapeNet for learned kernels.
+We compare Neural Kernel Fields (NKFs; Williams et al.
 (2022)) and Neural Kernel Surface Reconstruction (NKSR; Huang et al.
-(2023)).
+(2023)) against learnable Matérn 1/2 and 3/2.

### v3

@@
 where the function space $H:=H(\Omega)$ is usually taken to be a Reproducing Kernel Hilbert Space (RKHS) with associated reproducing kernel $k:\Omega\times\Omega\longrightarrow\mathbb{R}$, and $\lambda>0$.
-Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (Lorensen & Cline, 1987).
+Once $\hat{f}$ at hand, the reconstructed object’s surface is implicitly given as $\mathcal{\hat{S}}=\{x:\hat{f}(x)=0\}\subset\mathbb{R}^{d}$ and can be extracted using Marching Cubes (lorensen1987).
 It is quite easy to observe that, as $\lambda\rightarrow 0$, the optimization problem in Eq.
@@
 (2), while $\mathcal{\hat{S}}$ should interpolate the given points $\mathcal{X}$ exactly, the behavior in between and away from those points is solely determined by the induced norm $\|f\|_{H}$ of the function space, $H$, over which we are optimizing.
-The properties of $\|f\|_{H}$ are uniquely defined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) which states that solutions to Eq.
+The properties of $\|f\|_{H}$ are uniquely defined by the reproducing kernel, $k$; this is a direct consequence of the Representer Theorem (kimeldorf1970; schoelkopf2001) which states that solutions to Eq.
 (1) are of the form $f(x)=\sum_{i=1}^{m}\alpha_{i}k(x,x_{i})$, yielding $\|f\|_{H}^{2}=\alpha^{\top}K\alpha$.
@@
 Even further, if the chosen kernel has adjustable parameters, they can be used to adaptively (on a shape-by-shape basis) manipulate the inductive bias.
-Although already used in the mid-90s and early 2000s (Savchenko et al., 1995; Carr et al., 1997; 2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; Huang et al.
-(2023)) eventually evolving as the new state-of-the-art.
-While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (Williams et al., 2021; 2022) employ the first-order arc-cosine kernel.
-Introduced by Cho & Saul (2009), arc-cosine kernels have been shown to mimic the computation of (two-layer) fully-connected ReLU networks; indeed, if the layer-width tends to infinity, the first-order arc-cosine kernel is identified with the Neural Tangent Kernel (NTK; Jacot et al.
-(2018)) of the network.
+Although already used in the mid-90s and early 2000s (savchenko1995; car1997; car2001), only recently, kernel-based methods for 3D implicit surface reconstruction became extremely competitive, with Neural Kernel Surface Reconstruction (NKSR; huang2023) eventually evolving as the new state-of-the-art.
+While early works mostly focused on (polyharmonic) radial basis functions such as thin-plate splines, recent work (williams2021; williams2022) employ the first-order arc-cosine kernel.
+Introduced by cho2009, arc-cosine kernels have been shown to mimic the computation of two-layer fully-connected ReLU networks; indeed, in the infinite-width limit and with fixed bottom layer weights, the first-order arc-cosine kernel becomes the Neural Tangent Kernel (NTK; jacot2018) of the network.
 One important aspect of this kernel is that it does not have adjustable parameters.
@@
 Here, we show surface reconstructions from sparse point clouds of just 1,000 points.
-In this work, we suggest a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (Matérn, 1986; Stein, 1999), see Figures 1 and 2.
+In this work, we suggest a different family of kernels (with parameters $\nu,h>0$) for implicit surface reconstruction: Matérn kernels (matern1986; stein1999), see Figures 1 and 2.
 Contrary to arc-cosine kernels, Matérn kernels are stationary (hence translation invariant), spatially decaying (thus leading to sparse Gram matrices), and unify a variety of popular kernel functions, including the Laplace and Gaussian kernel.
@@
 From a practical perspective, we demonstrate that a simple change of the kernel function from arc-cosine to Matérn leads to a consistently improved reconstruction accuracy and a significant speedup.
-From a theoretical perspective, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (Tancik et al., 2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details).
-In addition, we prove that, as the layer-width approaches infinity, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activations; Sitzmann et al.
-(2020)) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs.
-Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace kernel (Matérn kernel with $\nu=1/2$) and the arc-cosine kernel are equal up to an affine transformation when restricted to the sphere.
+From a theoretical perspective, we argue that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings (tancik2020) enable coordinate-based Multi-layer Perceptrons (MLPs) to learn high-frequency functions in low-dimensional domains, effectively overcoming spectral bias (the slow convergence towards high-frequency details).
+In addition, we prove that, as the layer-width approaches infinity, with fixed bottom layer weights and under certain initializations, Matérn kernels are identified with the NTK of two-layer Siren (fully-connected MLPs with sine activations; sitzmann2020) networks—together with Fourier feature mappings the two arguably most influential methods to overcome spectral bias in coordinate-based MLPs.
+Lastly, we establish a connection between arc-cosine and Matérn kernels by showing that the Laplace (Matérn kernel with $\nu=1/2$) and first-order arc-cosine kernel are equal up to scaling and addition of another kernel function when restricted to the sphere.
 In summary, our key contributions are:
@@
 Moreover, we derive practical insights into how to choose the tunable bandwidth parameter based on a new bound of the $L_{2}$ reconstruction error.
-  • We propose data-dependent (i.e., learnable) Matérn kernels by leveraging the Neural Kernel Field (NKF) framework (Williams et al., 2022).
+  • We propose data-dependent (i.e., learnable) Matérn kernels by leveraging the Neural Kernel Field (NKF) framework (williams2022).
 Our experimental evaluation reveals that Matérn 1/2 and 3/2 are extremely competitive, outperforming the arc-cosine kernel while being significantly easier to implement (essentially two lines of standard PyTorch code), faster to compute, and scalable.
@@
 We briefly review some relevant literature about 3D implicit surface reconstruction from oriented point clouds, focusing on kernel-based methods.
-For a more in-depth overview, including neural-network-based reconstruction methods, see surveys by Berger et al.
-(2017); Huang et al.
-(2022b).
-Early kernel-based 3D surface reconstruction methods mostly employ Radial Basis Functions (RBFs) such as thin-plate splines (Savchenko et al., 1995), biharmonic RBFs (Carr et al., 1997; 2001), or Gaussian kernels (Schölkopf et al., 2005).
-Nowadays, the most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; Kazhdan & Hoppe (2013)) which, however, can itself be viewed as kernel method (Williams et al., 2021).
-Only recently, Neural Splines (NS; Williams et al.
-(2021)) proposed to use the so-called (first-order) arc-cosine kernel,
+For a more in-depth overview, including neural-network-based reconstruction methods, see surveys by berger2017; huang2022.
+Early kernel-based 3D surface reconstruction methods mostly employ Radial Basis Functions (RBFs) such as thin-plate splines (savchenko1995), biharmonic RBFs (car1997; car2001), or Gaussian kernels (schoelkopf2005).
+Nowadays, the most widely used surface reconstruction technique is Screened Poisson Surface Reconstruction (SPSR; kazhdan2013) which, however, can itself be viewed as kernel method (williams2021).
+Only recently, Neural Splines (NS; williams2021) proposed to use the so-called (first-order) arc-cosine kernel,
 $k_{\text{AC}}(x,y)=\frac{\|x\|\|y\|}{\pi}\left(\sin\theta+(\pi-\theta)\cos\theta\right),\quad\text{where}\quad\theta=\cos^{-1}\left(\frac{x^{\top}y}{\|x\|\|y\|}\right)$ (3)
 for implicit surface reconstruction, achieving state-of-the-art results that outperform classical surface reconstruction techniques and non-linear methods based on neural networks by a large margin.
-This method laid the cornerstone for Neural Kernel Fields (NKFs; Williams et al.
-(2022)) which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (Wilson et al., 2016).
-Based on SPSR, NeuralGalerkin (Huang et al., 2022a) proposed to learn basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense.
-Neural Kernel Surface Reconstruction (NKSR; Huang et al.
-(2023)) built upon NKF and NeuralGalerkin and proposed an all-purpose, highly scalable surface reconstruction method that is robust against noise and eventually became state-of-the-art.
+This method laid the cornerstone for Neural Kernel Fields (NKFs; williams2022) which attempts to make arc-cosine kernels learnable by passing input points through a task-specific neural network before evaluating the kernel function, similar to Deep Kernel Learning (wilson2016).
+Based on SPSR, NeuralGalerkin (huang2022_ng) proposed to learn basis functions inferred by a sparse convolutional network instead of using a fixed Bézier basis as in SPSR, hence can be seen as kernel method in the broader sense.
+Neural Kernel Surface Reconstruction (NKSR; huang2023) built upon NKF and NeuralGalerkin and proposed an all-purpose, highly scalable surface reconstruction method that is robust against noise and eventually became state-of-the-art.
 3 Matérn Kernels for Tunable Surface Reconstruction
-We propose to use the family of Matérn kernels (Matérn, 1986; Stein, 1999) for implicit surface reconstruction, being defined as
+We propose to use the family of Matérn kernels (matern1986; stein1999) for implicit surface reconstruction, being defined as
 $k_{\nu}(x,y)=\Phi_{\nu}(\|x-y\|)=\Phi_{\nu}(\tau)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}\tau}{h}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}\tau}{h}\right),$ (4)
@@
 $\Phi_{\infty}(\tau):=\lim_{\nu\rightarrow\infty}\Phi_{\nu}(\tau)=\exp\left(-\frac{\tau^{2}}{2h^{2}}\right).$ (8)
-For more information about Matérn kernels, please refer to Porcu et al. (2023). Next, we revisit some basic properties of Matérn kernels and place them in the context of surface reconstruction.
+For more information about Matérn kernels, please refer to porcu2023. Next, we revisit some basic properties of Matérn kernels and place them in the context of surface reconstruction.
 3.1 Basic Properties
@@
 Stationary.
-Matérn kernels are stationary, i.e., they only depend on the difference $x-y$ between two points $x,y\in\Omega$.
+Matérn kernels are stationary, i.e., they only depend on the difference xyx-y between two points $x,y\in\Omega$.
 If the distance is Euclidean, Matérn kernels are also isotropic; they only depend on $\|x-y\|$.
@@
 As opposed to arc-cosine kernels, Matérn kernels are spatially decaying; they tend to zero as the distance $\|x-y\|$ becomes large.
-Consequently, although Matérn kernels are technically not compact (or locally supported), because their value decays exponentially fast (Porcu et al., 2023), they can be considered effectively compact.
-From a computational perspective, this is a very desirable property as it leads to sparse Gram matrices when truncated with a special kernel (Genton, 2001), allowing the use of highly efficient and scalable sparse linear solvers.
+Consequently, although Matérn kernels are technically not compact (or locally supported), because their value decays exponentially fast (porcu2023), they can be considered effectively compact.
+From a computational perspective, this is a very desirable property as it leads to sparse Gram matrices when truncated with a special kernel (genton2001), allowing the use of highly efficient and scalable sparse linear solvers.
 Finally, we note that simply truncating any kernel does not yield a valid, positive definite kernel in general.
@@
 We begin by showing that Matérn kernels allow for tunable surface reconstruction in the same way as Fourier feature mappings help coordinate-based MLPs overcome spectral bias.
-As investigated in Tancik et al.
-(2020), a rapid decrease of the NTK’s eigenvalues implies an associated MLP’s slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components).
+As investigated in tancik2020, a rapid decrease of the NTK’s eigenvalues implies an associated MLP’s slow convergence to high-frequency components of the target function (up to the point where the network is practically unable to learn these components).
 Consequently, a smaller eigenvalue decay rate (EDR), i.e., a slower eigenvalue decay, leads to faster convergence to high-frequency content—in the context of implicit surface reconstruction, more detailed geometry.
-To overcome this slow convergence, a phenomenon known as spectral bias, Tancik et al.
-(2020) use a Fourier feature mapping of the form
+To overcome this slow convergence, a phenomenon known as spectral bias, tancik2020 use a Fourier feature mapping of the form
 $\gamma(x)=[a_{1}\cos(2\pi b_{1}^{\top}x),a_{1}\sin(2\pi b_{1}^{\top}x),\dots,a_{q}\cos(2\pi b_{q}^{\top}x),a_{q}\sin(2\pi b_{q}^{\top}x)]^{\top}$ (9)
@@
 To further study the dependence of the EDR on $\nu$ and $h$, we use:
-Theorem 1 (Seeger et al. (2008), Theorem 3).
+Theorem 1 (seeger2008, Theorem 3).
 The eigenvalues of Matérn kernels decay polynomially at rate
@@
 with bandwidth parameter $h>0$, and for finite smoothness $0<\nu<\infty$.
-Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by Li et al.
-(2024) for general input domains and distributions.
+Figure 3 visualizes the EDR for different values of $\nu$ and $h$ along with the EDR of the non-tunable, first-order arc-cosine kernel which equates to $\Theta(s^{-(1+d)/d})$ and has only been very recently proven by li2024 for general input domains and distributions.
 As seen, the smoother the kernel (i.e., the larger $\nu$) and the greater $h$, the faster the eigenvalues decay.
@@
 Surface reconstructions can be improved in practice by varying the bandwidth parameter, $h$, effectively tuning the kernels’ EDR (see also Figure 3).
-However, setting $h$ too small ($<0.5$) results in overfitting, while setting $h$ too big ($>10$) oversmooths (i.e., underfits) the true surface.
+However, setting $h$ too small (<0.5<0.5) results in overfitting, while setting $h$ too big (>10>10) oversmooths (i.e., underfits) the true surface.
 This is also reflected in the reconstruction error (measured using Chamfer distance), plotted as a function of $h$ on the right.
@@
 Moreover, $C_{d,\nu}:=(2^{d}\pi^{d/2}\Gamma(\nu+d/2)(2\nu)^{\nu})/\Gamma(\nu)$, and $\mathcal{F}[f]$ denotes the Fourier transform of a function $f$.
-For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., Santin & Schaback (2016).
+For more information on the bound, including details on $C_{\mathcal{X},\Omega}$, please see, e.g., santin2016.
 By inspecting Eq.
@@
 Next, we study the connection between Matérn kernels and neural networks.
-It is well known that the first-order arc-cosine kernel mimics the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., Cho & Saul (2009); Williams et al.
-(2021).
-Here, $[x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
-Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens (Sitzmann et al., 2020) if the layer-width approaches infinity. We claim the following:
+It is well known that the first-order arc-cosine kernel mimics the computation of a two-layer, infinite-width ReLU network $f:\mathbb{R}^{d}\longrightarrow\mathbb{R}$, $f(x)=m^{-1/2}\sum_{i=1}^{m}v_{i}[w^{\top}_{i}x+b_{i}]_{+}$ when the bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$ and biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$ are fixed from initialization and drawn from a standard normal distribution, see, e.g., cho2009; williams2021.
+Here, $x]_{+}:=\max\{0,x\}$ denotes the ReLU activation function.
+Under similar assumptions, we now show that Matérn kernels mimic the computation of two-layer Sirens (sitzmann2020) if the layer-width approaches infinity. We claim the following:
 Theorem 3.
@@
 Matérn kernels vs. arc-cosine kernels. Lastly, we aim to compare Matérn kernels with the previously used first-order arc-cosine kernel. Our analysis is based on:
-Theorem 4 (Chen & Xu (2021), Theorem 1).
+Theorem 4 (chen2021, Theorem 1).
 When restricted to the hypersphere $\mathbb{S}^{d-1}$, the RKHS of the Matérn kernel with smoothness $\nu=1/2$ (the Laplace kernel) include the same set of functions as the RKHS induced by the NTK of a fully-connected ($L\geq 2$)-layer ReLU network.
@@
 Corollary 5.
-When restricted to the hypersphere $\mathbb{S}^{d-1}$, (1) the RKHS of the Matérn kernel with smoothness $\nu=1/2$ (the Laplace kernel) is equal to the RKHS induced by the first-order arc-cosine kernel, implying that (2) the Laplace and arc-cosine kernel are equal up to an affine transformation.
+The RKHS of the Laplace kernel is equal to the RKHS induced by a sum of zeroth and first-order arc-cosine kernels, implying that, when restricted to $\mathbb{S}^{d-1}$, the Laplace and first-order arc-cosine kernel are equal up to constant scaling and addition of another kernel function.
 See Appendix E for a proof.
-For general input domains, our empirical results presented next support (at least for ($L=2$)-layer networks) the widely believed claim that the NTK is not significantly different from standard kernels (Belkin et al., 2018; Hui et al., 2019; Geifman et al., 2020), such as the Laplace kernel.
+For general input domains, our empirical results presented next support (at least for ($L=2$)-layer networks) the widely believed claim that the NTK is not significantly different from standard kernels (belkin2018; hui2019; geifman2020), such as the Laplace kernel.
 Indeed, we show that the Laplace kernel even outperforms the arc-cosine kernel.
 4 Experiments and Results
-We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (Chang et al., 2015) and the Surface Reconstruction Benchmark (SRB; Berger et al.
-(2013)) in Section 4.1.
-Moreover, Section 4.2 demonstrates Matérn kernels’ ability to reconstruct high-frequency textures on the Google Scanned Objects (GSO; Downs et al.
-(2022)) dataset and Objaverse (Deitke et al., 2023).
-Leveraging Neural Kernel Fields (Williams et al., 2022), we present and evaluate learnable Matérn kernels in Section 4.3.
+We systematically evaluated the effectiveness of Matérn kernels in the context of implicit surface reconstruction, presenting results on ShapeNet (chang2015) and the Surface Reconstruction Benchmark (SRB; berger2013) in Section 4.1.
+Moreover, Section 4.2 demonstrates Matérn kernels’ ability to reconstruct high-frequency textures on the Google Scanned Objects (GSO; downs2022) dataset and Objaverse (deitke2023).
+Leveraging Neural Kernel Fields (williams2022), we present and evaluate learnable Matérn kernels in Section 4.3.
 4.1 Surface Reconstruction on ShapeNet and SRB
 ShapeNet.
-We compare Matérn kernels in a sparse setting against the arc-cosine kernel on ShapeNet (using train/val/test split provided by Mescheder et al.
-(2019)).
+We compare Matérn kernels in a sparse setting against the arc-cosine kernel on ShapeNet (using train/val/test split provided by mescheder2019).
 To do so, we randomly sample $m=1,000$ surface points with corresponding normals for each shape.
@@
 F-Score $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 Matérn 1/2 93.6 93.7 93.7 4.05 4.02 4.02 Matérn 3/2 94.6 94.8 94.9 4.05 4.00 4.09 Matérn 5/2 92.4 93.8 92.9 6.42 5.65 6.91 Matérn $\infty$ 59.3 52.4 40.9 57.56 50.59 48.38 Arc-cosine 92.8 4.67 NS 90.6 4.74 SPSR 84.3 6.26    CD $\downarrow$ HD $\downarrow$ Matérn 1/2 0.21 4.43 Matérn 3/2 0.18 2.93 Matérn 5/2 0.58 22.52 Matérn $\infty$ 3.83 33.27 NS* 0.19 3.19 NS 0.17 2.85 SPSR 0.21 4.69 FFN 0.28 4.45 Siren 0.19 3.86 SAP 0.21 4.85 DiGS 0.18 3.55 VisCo 0.18 2.95 OG-INR 0.20 4.06
-[…]
+Table 1: Results on ShapeNet (left) and SRB (right) for non-learnable kernels.
+Arc-cosine kernel and Matérn $\nu\in\{\text{1/2},\text{3/2},\text{5/2},\infty\}$ are implemented in a unified framework and based on the exact same parameters.
+”NS*” denotes the best result we could achieve by running the official implementation of NS, see Appendix G.
+Bold marks best result, underline second best.
 Quantitative results in terms of F-Score (with a threshold of 0.01) and Chamfer distance (CD; always reported $\times$ 103) can be found in Table 1.
-Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (Kazhdan & Hoppe, 2013).
+Matérn 1/2 and 3/2 consistently outperform the arc-cosine kernel as well as popular SPSR (kazhdan2013).
 In addition, our experiments show that the bandwidth parameter, $h$, can be used to tune the surface reconstruction (lower the error), while not being too sensitive in practice.
@@
 For this, we implemented a highly scalable version of Matérn kernels, closely following the NS framework to solve the KRR problem in Eq.
-(1) using Nyström sampling (with 15,000 samples) and the FALKON (Rudi et al., 2017) solver.
-We compare against NS and SPSR, as well as kernel-free methods, Fourier Feature Networks (FFNs; Tancik et al.
-(2020)), Siren, SAP (Peng et al., 2021), DiGS (Ben-Shabat et al., 2022), VisCo (Pumarola et al., 2022), and OG-INR (Koneputugodage et al., 2023).
+(1) using Nyström sampling (with 15,000 samples) and the FALKON (rudi2017) solver.
+We compare against NS and SPSR, as well as kernel-free methods, Fourier Feature Networks (FFNs; tancik2020), Siren, SAP (peng2021), DiGS (shabat2022), VisCo (pumarola2022), and OG-INR (koneputugodage2023).
 All methods are optimization-based (FFN and Siren are overfitted to a single shape) and use surface normals.
@@
 Notably, we did not tune Matérn kernels’ bandwidth parameter for this experiment.
-We also show qualitative results on the challenging Objaverse (Deitke et al., 2023) dataset in Figure 7.
+We also show qualitative results on the challenging Objaverse (deitke2023) dataset in Figure 7.
 We chose Objaverse as it includes extremely high-resolution and complex textures.
@@
 4.3 Data-dependent Matérn Kernels
-We leverage the Neural Kernel Field (NKF) framework introduced by Williams et al.
-(2022) to make Matérn kernels learnable.
+We leverage the Neural Kernel Field (NKF) framework introduced by williams2022 to make Matérn kernels learnable.
 Specifically, NKF proposed to feed points $x,y\in\mathbb{R}^{d}$ through an input-conditioned neural network $\varphi:\mathbb{R}^{d}\longrightarrow\mathbb{R}^{q}$ before evaluating a kernel function, $k^{\varphi}(x,y;\mathcal{D})=k\left([x,\varphi(x;\mathcal{D})],[y,\varphi(y;\mathcal{D})]\right)$.
@@
 We set $q=32$ for our experiments.
-Please see Williams et al.
-(2022) for details.
+Please see williams2022 for details.
 Sparse surface reconstruction and extreme generalization.
-We compare learned Matérn kernels against the original NKF framework (with the arc-cosine kernel) and NKSR (Huang et al., 2023) in a sparse setting on ShapeNet.
+We compare learned Matérn kernels against the original NKF framework (with the arc-cosine kernel) and NKSR (huang2023) in a sparse setting on ShapeNet.
 Again, we sample $m=1,000$ surface points and corresponding normals for each shape and set $h=1$ for all Matérn kernels.
@@
 IoU $\uparrow$ CD $\downarrow$ NC $\uparrow$ Matérn 1/2 95.3 2.58 95.6 Matérn 3/2 94.9 2.70 95.3 Matérn 5/2 93.3 3.07 94.9 Matérn $\infty$ 92.1 3.39 94.2 NKF* 94.7 2.70 95.2 NKF 94.7 2.65 95.4 NKSR 95.6 2.34 95.4     Train only on chairs, test on all IoU $\uparrow$ CD $\downarrow$ NC $\uparrow$ Time/epoch Matérn 1/2 93.4 3.06 94.3 7.71 min Matérn 3/2 92.8 3.34 94.2 8.26 min Matérn 5/2 90.5 4.11 93.9 8.56 min Matérn $\infty$ 84.7 6.70 92.6 7.74 min NKF* 92.8 3.30 94.1 31.98 min NKSR 89.6 2.70 94.1 41.66 min
-[…]
+Table 4: Results on ShapeNet for learned kernels.
+”NKF*” denotes a re-implemented variant of NKF.
+NKF* and Matérn $\nu\in\{\text{1/2},\text{3/2},\text{5/2},\infty\}$ are based on the same framework and parameters; they differ only in the employed kernel.
+Runtime is measured on a single NVIDIA A100 with a batch size of one to ensure fair comparison.
+Bold marks best result, underline second best.
 No noise ($\sigma=0$) Small noise ($\sigma=0.0025$) Big noise ($\sigma=0.005$) IoU $\uparrow$ CD $\downarrow$ IoU $\uparrow$ CD $\downarrow$ IoU $\uparrow$ CD $\downarrow$ 0.5 1 2 0.5 1 2 0.5 1 2 0.5 1 2 0.5 1 2 0.5 1 2 Matérn 1/2 93.3 93.6 93.5 2.88 2.85 2.87 91.9 92.1 92.1 3.03 3.04 3.07 89.5 89.6 89.6 3.37 3.36 3.39 Matérn 3/2 92.7 92.7 92.3 3.08 3.20 3.25 89.7 89.8 89.7 3.48 3.69 3.54 85.8 85.5 86.3 4.23 4.32 4.15 Matérn 5/2 91.8 90.5 88.8 3.58 3.69 4.18 87.9 87.6 87.0 4.29 4.48 4.43 83.4 82.2 84.3 4.90 5.03 5.11 Matérn $\infty$ 86.4 85.8 86.8 5.80 5.07 4.75 81.3 83.3 81.0 7.27 5.62 6.45 76.7 80.7 79.3 8.27 5.90 6.67 NKF* 93.2 2.97 92.0 3.12 89.5 3.39 NKSR 91.1 2.65 90.1 2.98 88.4 3.41
@@
 Appendix A On the Stationarity of Matérn Kernels
-As noted in Section 3.1 of the main paper, Matérn kernels are stationary: they only depend on the difference $x-y$ between two points $x,y\in\Omega$, hence being translation invariant.
-This is because $k(x-y)=k(x+t-(y+t))$ for $t\in\mathbb{R}^{d}$ and all stationary kernels, $k$.
+As noted in Section 3.1 of the main paper, Matérn kernels are stationary: they only depend on the difference xyx-y between two points $x,y\in\Omega$, hence being translation invariant.
+This is because k(xy)=k(x+t(y+t))k(x-y)=k(x+t-(y+t)) for $t\in\mathbb{R}^{d}$ and all stationary kernels, $k$.
 In contrast, the arc-cosine kernel is not translation invariant as its value depends on the absolute positions, $x,y$, see Eq.
-(3) of the main paper (i.e., the arc-cosine kernel can not be written as a function of $x-y$).
+(3) of the main paper (i.e., the arc-cosine kernel can not be written as a function of xyx-y).
 In fact, the arc-cosine kernel is non-stationary.
@@
 $k(\tau)=k(x-y)=\int_{\mathbb{R}^{d}}e^{i\omega^{\top}(x-y)}\mathrm{d}\mu(\omega)=\mathbb{E}_{\omega\sim\mu}\left[\cos(\omega^{\top}(x-y))\right]$
-because $k$ is real and for even spectral densities. On the other hand, Rahimi & Recht (2007) showed
+because $k$ is real and for even spectral densities. On the other hand, rahimi2007 showed
 $k(\tau)=k(x-y)$ $=\mathbb{E}_{\omega\sim\mu}[\cos(\omega^{\top}(x-y))]$
@@
 $=\mathbb{E}_{\omega\sim\mu,b\sim\mathcal{U}(0,2\pi)}[2\cos(\omega^{\top}x+b)\cos(\omega^{\top}y+b)],$
-where ()(*)( ∗ ) follows because $b$ is uniformly distributed in $[0,2\pi]$ (hence, the inner expectation w.r.t. $b$ becomes zero).
+where ()(*) follows because $b$ is uniformly distributed in $0,2\pi]$ (hence, the inner expectation w.r.t. $b$ becomes zero).
 We will now prove Theorem 3, copied below for convenience:
@@
 $f(x)=\sqrt{\frac{2}{m}}\sum_{i=1}^{m}v_{i}\sin(w_{i}^{\top}x+b_{i})$
-with bottom-layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$, biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$, and top-layer weights $v=(v_{1},v_{2},\dots,v_{m})\in\mathbb{R}^{m}$.
+with bottom layer weights $W=(w_{1},w_{2},\dots,w_{m})\in\mathbb{R}^{d\times m}$, biases $b=(b_{1},b_{2},\dots,b_{m})\in\mathbb{R}^{m}$, and top-layer weights $v=(v_{1},v_{2},\dots,v_{m})\in\mathbb{R}^{m}$.
 Assume that $W$ and $b$ are fixed (that is, we only allow the second layer to be trained), and initialized according to some distribution.
@@
 $k_{\text{NTK}}(x,y)$ $=\mathbb{E}_{(w,b)}\left[2\cos(w^{\top}x+b)\cos(w^{\top}y+b)\right]-2\underbrace{\mathbb{E}_{(w,b)}\left[\cos(w^{\top}(x+y)+2b)\right]}_{=0\,(*)}$
-${=}k_{\nu}(x,y).$
-Here, ()(*)( ∗ ) follows because $b$ is uniformly distributed, and ()(**)( ∗ ∗ ) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see, e.g., Eq.
-(10) in Kanagawa et al.
-(2018).
+$\hskip-10.27777pt\overset{(**)}{=}k_{\nu}(x,y).$
+Here, ()(*) follows because $b$ is uniformly distributed, and ()(**) since $w$ is distributed according to $p_{\nu}$ which is, in fact, the spectral density of $k_{\nu}$, see, e.g., Eq.
+(10) in kanagawa2018.
 ∎
@@
 Remark 7.
-The NTK of a Siren can be associated with any stationary kernel if the bottom-layer weights in Eq. (13) are initialized according to the corresponding kernel’s spectral density.
+The NTK of a Siren can be associated with any stationary kernel if the bottom layer weights in Eq. (13) are initialized according to the corresponding kernel’s spectral density.
 This is a direct consequence of Bochner’s theorem.
@@
 We first re-state the following result, needed to prove the second part of Corollary 5:
-Theorem 8 (Saitoh & Sawano (2016), Theorem 2.17).
+Theorem 8 (saitoh2016, Theorem 2.17).
 Let $k_{1}$ and $k_{2}$ be two positive definite kernels. Denote by $H_{1},H_{2}$ the RKHSs induced by $k_{1}$ and $k_{2}$. Then,
@@
 Proof.
-(1) Follows from Theorem 4 by noting that the NTK of a fully-connected $(L=2$)-layer ReLU network is the first-order arc-cosine kernel, see, e.g.Cho & Saul (2009); Williams et al.
-(2021).
-For (2), we know that $H_{\nu}\subset H_{\text{AC}}$ and $H_{\text{AC}}\subset H_{\nu}$, so, based on Theorem 8, $k:=\gamma^{2}k_{\text{AC}}-k_{\nu}$ and $\bar{k}:=\bar{\gamma}^{2}k_{\nu}-k_{\text{AC}}$ are both valid kernels for $\gamma,\bar{\gamma}>0$.
-Re-arranging both equations, we have
-$k_{\text{AC}}=ak_{\nu}+b,\quad\text{where}\quad a=1/\gamma^{2},b=k/\gamma^{2}$
-and, equivalently,
-$k_{\nu}=\bar{a}k_{\text{AC}}+\bar{b},\quad\text{where}\quad\bar{a}=1/\bar{\gamma}^{2},\bar{b}=\bar{k}/\bar{\gamma}^{2}.$
-This shows that the Laplace and arc-cosine kernel are indeed related by an affine transformation when restricted to $\mathbb{S}^{d-1}$, concluding the proof. ∎
+The first part follows from Theorem 4 by observing that the NTK of a fully-connected $(L=2$)-layer ReLU network (with all parameters being trained) is
+$k_{\widetilde{\text{AC}}}(x,y)=\underbrace{k_{\text{AC}}(x,y)}_{\text{first-order}}+x^{\top}y\underbrace{\frac{1}{\pi}(\pi-\theta)}_{\text{zeroth-order}}$ (14)
+see, e.g.chen2021.
+The second part follows from Theorem 4, noting that $H_{\widetilde{\text{AC}}}\subset H_{\text{1/2}}$ when restricted to $\mathbb{S}^{d-1}$; so, based on Theorem 8, $k:=\gamma^{2}k_{\text{1/2}}-k_{\widetilde{\text{AC}}}$ is a valid kernel for $\gamma>0$.
+Re-arranging the equation, we have
+$k_{\text{1/2}}=ak_{\widetilde{\text{AC}}}+b,\quad\text{where}\quad a=1/\gamma^{2},b=k/\gamma^{2}.$ (15)
+Now, denote by $k_{\text{AC}}^{(0)}(x,y)$ the zeroth-order arc-cosine kernel as defined in Eq.
+(14), and further set $\hat{k}_{\text{AC}}^{(0)}(x,y)=x^{\top}yk_{\text{AC}}^{(0)}(x,y)$.
+The latter is a product between two kernel functions, hence a valid, i.e., positive definite, kernel itself.
+Plugin Eq.
+(14) into Eq.
+(15) yields
+$\begin{split}k_{\text{1/2}}&=a(k_{\text{AC}}+\hat{k}_{\text{AC}}^{(0)})+b\\ &=ak_{\text{AC}}+c\end{split}\quad\text{with}\quad c=a\hat{k}_{\text{AC}}^{(0)}+b,$ (16)
+where $c$ is again a valid kernel function since it is a sum of two kernels.
+This shows that the Laplace and first-order arc-cosine kernel are indeed related by a constant scaling factor $a$ and the addition of a kernel function $c$ when restricted to $\mathbb{S}^{d-1}$, concluding the proof.
+∎
+In other words, Corollary 5 states that, when restricted to the sphere, the Laplace kernel is a linear combination of two kernel functions: the first-order arc-cosine kernel and another, valid kernel function (that accounts for the residuals).
+This holds also in the other direction.
 Appendix F Additional Details for ShapeNet Experiment
@@
 F.1 Implementation Details
-Following common practice (see, e.g., Williams et al.
-(2022)), we use finite differences to approximate the gradient part of the KRR problem in Eq.
+Following common practice (see, e.g., williams2022), we use finite differences to approximate the gradient part of the KRR problem in Eq.
 (1) of the main paper.
 Specifically, denote $\mathcal{X}^{\prime}:=\mathcal{X}^{+}\cup\mathcal{X}^{-}$, where $\mathcal{X}^{+}:=\{x_{1}+{\epsilon}n_{1},x_{2}+{\epsilon}n_{2},\dots,x_{m}+{\epsilon}n_{m}\}$ and $\mathcal{X}^{-}:=\{x_{1}-{\epsilon}n_{1},x_{2}-{\epsilon}n_{2},\dots,x_{m}-{\epsilon}n_{m}\}$ for a fixed ${\epsilon}>0$.
-The Representer Theorem (Kimeldorf & Wahba, 1970; Schölkopf et al., 2001) tells us that the solution to Eq.
+The Representer Theorem (kimeldorf1970; schoelkopf2001) tells us that the solution to Eq.
 (1) is of the form
@@
 Appendix G Further Details on the Surface Reconstruction Benchmark
-This section details the experimental setting that has been used to run Neural Splines (NS; Williams et al.
-(2021)) on SRB in Section 4.1 of the main paper.
+This section details the experimental setting that has been used to run Neural Splines (NS; williams2021) on SRB in Section 4.1 of the main paper.
 Moreover, we present per-object metrics and additional qualitative results.
@@
 We run on four different GPUs, including NVIDIA RTX A2000, A40, A100, and V100 (which has been used by the authors of Neural Splines), and tested different configurations of the following parameters (see code): grid_size, eps, cg-max-iters, cg-stop-thresh, outer-layer-variance.
-In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; Williams et al.
-(2021) reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively.
+In Table 1 of the main paper, we report the best result we could achieve on a V100, which is 0.19 for Chamfer and 3.19 for the Hausdorff distance; williams2021 reported 0.17 and 2.85 for Chamfer and Hausdorff distance, respectively.
 In order to achieve this, we had to lower outer-layer-variance form $10^{-3}$ per default to $10^{-5}$.
@@
 Due to the absence of further information about the experimental setting in the original paper, we used default values provided in the repository for the rest of the parameters.
-Runtime was measured on a single NVIDIA V100, similar to Williams et al.
-(2021).
+Runtime was measured on a single NVIDIA V100, similar to williams2021.
 G.2 Additional Results
@@
 Bold marks best result, underline second best.
-Figure 10: Additional qualitative results on SRB for non-learnable kernels. We compare Neural Splines (NS; Williams et al. (2021)) to Matérn 1/2 and 3/2.
+Figure 10: Additional qualitative results on SRB for non-learnable kernels. We compare Neural Splines (NS; williams2021) to Matérn 1/2 and 3/2.
 Appendix H Additional Details For Texture Reconstruction Experiment
@@
 To extract a textured surface mesh from the predicted four-dimensional volume, we use a two-stage process.
-First, we apply Marching Cubes (Lorensen & Cline, 1987) to $f^{\prime}_{1}$, and then trilinearily interpolate RGB colors at the previously predicted surface points.
+First, we apply Marching Cubes (lorensen1987) to $f^{\prime}_{1}$, and then trilinearily interpolate RGB colors at the previously predicted surface points.
 H.2 Evaluation Protocol
-Our quantitative evaluation on the Google Scanned Objects (GSO; Downs et al.
-(2022)) dataset closely follows Mitchel et al.
-(2024), in which authors compute image-based evaluation metrics on the texture atlas to quantify reconstruction ability.
+Our quantitative evaluation on the Google Scanned Objects (GSO; downs2022) dataset closely follows mitchel2024, in which authors compute image-based evaluation metrics on the texture atlas to quantify reconstruction ability.
 Specifically, we first sample $m$ surface points along with corresponding normals and per-point RGB colors from shapes in GSO.
@@
 Figure 11: Additional qualitative results on ShapeNet for learned kernels.
-We compare Neural Kernel Fields (NKFs; Williams et al.
-(2022)) and Neural Kernel Surface Reconstruction (NKSR; Huang et al.
-(2023)) against learnable Matérn 1/2 and 3/2.
+We compare Neural Kernel Fields (NKFs; williams2022) and Neural Kernel Surface Reconstruction (NKSR; huang2023) against learnable Matérn 1/2 and 3/2.